### v4

# Geometry, Computation, and Optimality in Stochastic Optimization

Abstract
We study computational and statistical consequences of problem geometry in stochastic and online optimization. By focusing on constraint set and gradient geometry, we characterize the problem families for which stochastic- and adaptive-gradient methods are (minimax) optimal and, conversely, when nonlinear updates—such as those mirror descent employs—are necessary for optimal convergence. When the constraint set is quadratically convex, diagonally pre-conditioned stochastic gradient methods are minimax optimal. We provide quantitative converses showing that the “distance” of the underlying constraints from quadratic convexity determines the sub-optimality of subgradient methods. These results apply, for example, to any $\ell_{p}$-ball for $p<2$, and the computation/accuracy tradeoffs they demonstrate exhibit a striking analogy to those in Gaussian sequence models.
1 Introduction
The default procedures for solving the stochastic optimization problem
[…]
where $\{F(\cdot,s),s\in\mathcal{S}\}$ are convex functions $F(\cdot,s):\mathbf{R}^{n}\to\mathbf{R}$, $P$ is a distribution on $\mathcal{S}$, and $X\subset\mathbf{R}^{n}$ is a closed convex set, are variants of the stochastic subgradient method, where one iteratively draws $S_{k}\stackrel{{\scriptstyle\textup{iid}}}{{\sim}}P$ and updates
$x_{k+1}:=x_{k}-\alpha_{k}g_{k},~{}~{}\mbox{where}~{}~{}g_{k}\in\partial F(x_{k},S_{k}).$ (2)
The simplicity and scalability of this update make stochastic subgradient methods the de facto choice for large-scale optimization [33, 28, 8]. The geometry of the underlying underlying constraint set $X$ and subgradients $\partial F(\cdot,s)$ impact the performance of algorithms for problem (1), and so a question arises: are such linear updates (2) enough to obtain (minimax rate) optimal convergence guarantees for the problem (1), or does does the structure of the problem necessitate nonlinearity to achieve optimization efficiency? Convergence guarantees for stochastic gradient methods depend on the $\ell_{2}$-diameter of $X$ and $\partial F(\cdot,s)$, while for non-Euclidean geometries (e.g. when $X$ is an $\ell_{1}$- or $\ell_{\infty}$-ball) mirror descent, dual averaging and adaptive gradient methods provide better convergence guarantees [27, 28, 5, 29, 19, 14]. We investigate these gaps by precisely quantifying convergence for different method families, highlighting a particular way to trade between computational power—which we treat as whether purely linear operations suffice to optimally solve problem (1), or nonlinear updates are necessary—and optimization and statistical efficiency.
To set the stage, let us revisit Donoho, Liu, and MacGibbon’s study of optimal estimation in Gaussian sequence models [17]. One observes a vector $x\in X$ corrupted by Gaussian noise,
$y=x+\mathsf{N}(0,\sigma^{2}I),$ (3)
and seeks to estimate $x$. For such problems, one can consider linear estimators—$\widehat{x}=AY$ for a $A\in\mathbf{R}^{n\times n}$—or potentially non-linear estimators
$\widehat{x}=\Phi(y)$
where $\Phi:\mathbf{R}^{n}\to X$ is otherwise arbitrary. When $X$ is quadratically convex, meaning the set $X^{2}:=\{(x_{j}^{2})_{j\geq 1}\mid x\in X\}$ is convex, Donoho et al. show there exist minimax rate-optimal linear estimators; conversely, there are non-quadratically convex $X$ for which rate-optimal estimators $\widehat{x}$ must be nonlinear in $y$. In particular, as we discuss in Section 5, this gap depends on the difference between the Kolmogorov (linear) $n$-width of $X$ and its “nonlinear” $n$-width, that is,
$0pt^{2}(n):=\sup_{v\in\textup{Conv}(X^{2})}\sum_{j>n}v_{(j)}~{}~{}~{}\mbox{versus}~{}~{}~{}w_{\textup{nl}}^{2}(n):=\sup_{v\in X^{2}}\sum_{j>n}v_{(j)},$ (4)
where $|v_{(1)}|\geq|v_{(2)}|\geq\cdots$ denote the elements of $v$ sorted by magnitude. We show how these results follow from convex duality, and the difference between $0pt^{2}(n)$ and $w_{\textup{nl}}^{2}(n)$ allows a quantitative characterization of how far $X$ is from being quadratically convex and the impact this distance has on the (sub)optimality of linear estimators.
Our results show how stochastic and online convex optimization analogize these sequence models. To build the analogy, consider dual averaging [29], where for a strongly convex $h:X\to\mathbf{R}\cup\{+\infty\}$, one iteratively receives $S_{k}\in\mathcal{S}$, chooses $g_{k}\in\partial F(x_{k},S_{k})$, and for a stepsize $\alpha_{k}>0$ updates
$x_{k+1}:=\mathop{\textup{argmin}}_{x}\bigg{\{}\sum_{i\leq k}g_{i}^{\top}x+\frac{1}{\alpha_{k}}h(x)\bigg{\}},$ (5)
By taking $h(x)=h_{0}(x)+\mathbf{1}_{X}(x)$, where $h_{0}$ is strongly convex on $X$ and $\mathbf{1}_{X}(x)=0$ when $x\in X$ and +$\infty$+ ∞ otherwise, $h$ can incorporate constraints in the update (5). When $h$ is Euclidean, that is, $h(x)=\frac{1}{2}x^{\top}Ax$ for some $A\succ 0$, the updates are linear in the observed gradients $g_{i}$, as
$x_{k}=-\alpha_{k}A^{-1}\sum_{i\leq k}g_{i}.$
Adapting common practice in online learning [11] (and in sequence model estimation (3)), we allow improper methods, which may output estimates $\widehat{x}\not\in X$, saying that $\widehat{x}$ is $\epsilon$-sub-optimal relative to $X$ if
$f(\widehat{x})\leq\inf_{x\in X}f(x)+\epsilon.$
Drawing a parallel between $\Phi$ in the Gaussian sequence model and $h$ in dual averaging (5), we show that because of duality gaps in certain min-max problems, a dichotomy holds for stochastic and online convex optimization similar to that holding for the Gaussian sequence model: if $X$ is quadratically convex, there is a Euclidean $h$ (yielding linear updates (5)) that is minimax rate optimal for problem (1), while there exist non-quadratically convex $X$ for which Euclidean distance-generating $h$ are arbitrarily suboptimal. Taking a computational perspective, this means that for some problems one must use more sophisticated methods than “linear” updates. We show that this analogy holds, though the measurement of a set’s deviance from quadratic convexity, and hence the gap in attainable performance between linear and nonlinear methods, differs between Gaussian sequence models and stochastic optimization: there are constraint sets $X$ for which linear estimators are (rate) optimal in the Gaussian sequence model but not for stochastic optimization, and vice versa.
More precisely, we prove that for orthosymmetric quadratically convex bodies $X$, subgradient methods with a fixed diagonal re-scaling are minimax rate optimal. This guarantees that for a large collection of constraints (e.g. $\ell_{2}$-balls, weighted $\ell_{p}$-bodies for $p\geq 2$, or hyperrectangles) a diagonal re-scaling suffices. This is important in, e.g., machine learning problems of appropriate geometry, such as in linear classification problems where the data (features) are sparse, so using a dense predictor $x$ is natural [19, 20]. Conversely, we show that if the constraint set $X$ is a (scaled) $\ell_{p}$-ball, $21\leq p<$ 2, then, considering unconstrained updates (5), the regret of the best method of linear type (i.e. $h$ quadratic) can be $\sqrt{n/\log n}$ times larger than the minimax rate in online convex optimization. As part of this, we provide new information-theoretic lower bounds on optimization for general convex constraints $X$. In contrast to the frequent (but illogical) practice of comparing convergence upper bounds, we demonstrate that the gap between linear and non-linear methods must hold. Sections 4.2 and 5 also show how the departure from quadratic convexity affects convergence guarantees: comparing the $\ell_{1}$ diameters of $X$ and its second-order lifts via
$\sup_{x\in X}\left\|{x}\right\|_{1}~{}~{}\mbox{versus}~{}~{}\sup_{v}\left\{\left\|{v}\right\|_{1}\mid v^{2}\in\textup{Conv}\{\textup{diag}(xx^{\top}),x\in X\})\right\},$
the gap between the left and right quantities (essentially) characterizes the gap in performance between linear and nonlinear methods for stochastic optimization, while Kolmogorov $n$-widths (4) capture that in the Gaussian sequence model.
We extend our results to an additional computational consideration: whether an algorithm must be adaptive, that is, it must change its update rules over time based on observations. We demonstrate that non-adaptive linear methods necessarily suffer slower convergence rates than adaptive methods in online problems. One perspective on our results is thus computational, though with a different angle than most current work on tradeoffs between statistics and computational complexity. Much of this literature takes as inspiration the classical perspective that the gap between polynomial and non-polynomial time algorithms forms the great watershed in computational complexity, thus necessitating a class of “hard” problems while allowing essentialy arbitrary algorithms [6, 9, 10]. We take an alternative perspective that allows more nuance in the types of convergence rates we can achieve—differentiating between various polynomials—by restricting the algorithms we consider to those in families common in optimization.
Our conclusions relate to the growing literature in adaptive algorithms [4, 19, 30, 31, 15]. Our results effectively prescribe that these adaptive algorithms are useful when the constraint set is quadratically convex, as this guarantees a minimax optimal diagonal pre-conditioner. More, different sets suggest different regularizers. For example, when the constraint set is a hyperrectangle, AdaGrad has regret at most $\sqrt{2}$ times that of the best post-hoc pre-conditioner, which we show is minimax optimal, while (non-adaptive) standard gradient methods can be $\sqrt{n}$ suboptimal on such problems. Conversely, our results show that these diagonal rescaling methods may be quite rate sub-optimal for non-quadratically convex constraint sets. Our results thus clarify existing convergence guarantees [29, 28, 19, 43]: when the geometry of $X$ and $\partial F$ is appropriate for adaptive gradient methods or Euclidean algorithms, one should use them; when it is not—the constraints $X$ are not quadratically convex—other methods admit substantially better convergence.
Notation
[…]
2 Preliminaries and Background
We begin by reviewing the classical results in Gaussian sequence models and presenting and defining the minimax framework in which we analyze procedures. We also review standard stochastic subgradient methods and introduce the relevant geometric notions of convexity we require. As part of this, we give a new argument showing the optimality of linear estimators for Gaussian sequence models when the underlying constraint set is quadratically convex (which we define presently).
Quadratic convexity and orthosymmetry
A few geometric quantities are central to our development. For a set $X$, let $X^{2}:=\{x^{2},x\in X\}$ denote its (elementwise) square. The set $X$ is quadratically convex if $X^{2}$ is convex; typical examples of quadratically convex sets are weighted $\ell_{p}$ bodies for $p\geq 2$ or hyperrectangles. We let $\mathsf{QHull}(X)$ be the quadratic convex hull of $X$, meaning the smallest convex and quadratically convex set containing $X$. The set $X\subset\mathbf{R}^{n}$ or $X\subset\mathbf{R}^{\mathbf{N}}$ is orthosymmetric if it is invariant to flipping the signs of any coordinate: if $x\in X$ then $\sigma_{j}\in\{\pm 1\}$ implies $(\sigma_{j}x_{j})_{j\geq 1}\in X$. Similarly, a norm $\gamma$ is orthosymmetric if $\gamma(g)=\gamma(|g|)$ for all $g$, and $\gamma$ is quadratically convex if it induces a quadratically convex unit ball $\mathbf{B}_{\gamma}(0,1)$. For any set $X$, we define the squared convex hull and square root
$\mathsf{SqHull}(X):=\textup{Conv}\left\{(x_{j}^{2})\mid x\in X\right\}~{}~{}\mbox{and}~{}~{}\sqrt{\mathsf{SqHull}(X)}=\{(\sqrt{y_{j}})\mid y\in\mathsf{SqHull}(X)\},$
the latter of which is always convex by the concavity of the square root. For orthosymmetric $X$,
$\mathsf{QHull}(X)=\left\{s\odot x\mid x\in\sqrt{\mathsf{SqHull}(X)},s_{j}\in\{\pm 1\}~{}\mbox{for~{}all~{}}j\right\}.$
2.1 The Gaussian sequence model
Gaussian sequences provide a model for analyzing parametric and nonparametric statistical procedures, and tools developed in their analysis form a bedrock of modern statistical estimation [40, 25]; we provide some perspective on estimation in the sequence model. In the Gaussian sequence model, we begin with a (typically convex and compact) set $X\subset\mathbf{R}^{n}$ or in sequence space $\mathbf{R}^{\mathbf{N}}$, and for an unknown $x\in X$ observe $y=x+\xi$, where $\xi\sim\mathsf{N}(0,\sigma^{2}I)$. The goal is to estimate $x$ in some sense optimally, and frequently one considers sequences with $X\subset\mathbf{R}^{n}$ and $\sigma^{2}$ scaling as $1/n$, which analogizes estimation with $n$ observations, so that rates of convergence as $\sigma^{2}\downarrow 0$ become the main focus [25]. An interesting point of contrast is when linear estimators are sufficient to achieve (near) optimal performance or nonlinear estimators are necessary. An estimator $\widehat{x}=\widehat{x}(y)$ is linear if it has the form $\widehat{x}=Ay$ for a linear operator $A$ and nonlinear otherwise. We consider the risk
$R(\widehat{x},x):=\mathbf{E}\left[\|{\widehat{x}(y)-x}\|_{2}^{2}\right],$
where for linear estimators of the form $\widehat{x}=Ay$, we use the shorthand
$R(A,x)=\mathbf{E}\left[\left\|{Ay-x}\right\|_{2}^{2}\right]=\mathbf{E}\left[\left\|{(A-I)x+A\xi}\right\|_{2}^{2}\right]=\left\|{(A-I)x}\right\|_{2}^{2}+\sigma^{2}\left\|{A}\right\|_{\rm Fr}^{2}.$ (6)
The maximum risk of an estimator over the set $X$ is
$R^{*}(\widehat{x},X):=\sup_{x\in X}R(\widehat{x},x),$
while the minimax risk and linear minimax risk are
$R^{*}(X):=\inf_{\widehat{x}}R^{*}(\widehat{x},X)~{}~{}~{}\mbox{and}~{}~{}~{}R_{\textup{lin}}^{*}(X):=\inf_{A}R^{*}(A,X).$
Donoho et al. [17] prove fundamental results relating the minimax risk and linear minimax risk for the Gaussian sequence model, and among their main results is that if the set $X$ is an orthosymmetric quadratically convex body, then $R_{\textup{lin}}^{*}(X)\leq 1.25R^{*}(X)$, and moreover, $R_{\textup{lin}}^{*}(X)=\sup_{H\subset X}R_{\textup{lin}}^{*}(H)$, where $H$ is a (hyper)rectangle. We provide an alternative approach to some of these arguments via convex duality, which allows us to put these arguments and the rest of our development on similar intellectual footing, though we defer proofs to appendices. In brief, the Sion and Fan minimax theorems [37, 22], coupled with quadratic lifts of the set $X$, play an essential role in all of our results.
Proposition 1.
Assume $\sigma^{2}>0$ and $X\subset\mathbf{R}^{n}$ is an orthosymmetric convex body. Then the matrix $A$ minimizing $R^{*}(A,X)$ is diagonal and unique. Moreover,
$\inf_{A}R^{*}(A,X)=\inf_{d\in\mathbf{R}^{n}}\sup_{x\in X}\left\{\sum_{j=1}^{n}(d_{j}-1)^{2}x_{j}^{2}+\sigma^{2}d_{j}^{2}\right\}.$
See Appendix A.1 for the proof of the result. Then a quick proof via saddle point duality characterizes the linear minimax risk for the Gaussian sequence model (see Appendix A.2).
Corollary 1.
Let $X\subset\mathbf{R}^{n}$ be an orthosymmetric quadratically convex body. Then
$\inf_{A}R^{*}(A,X)=\sup_{q\in\mathsf{QHull}(X)}\sum_{j=1}^{n}\frac{\sigma^{2}q_{j}^{2}}{q_{j}^{2}+\sigma^{2}}.$
An approximation argument extends Corollary 1 to sequence space (see Appendix A.3).
Corollary 2.
Let $X\subset\mathbf{R}^{\mathbf{N}}$ be orthosymmetric and compact for $\ell_{2}(\mathbf{N})$. Then
$\inf_{A}R^{*}(A,X)=\sup_{q\in\mathsf{QHull}(X)}\sum_{j=1}^{\infty}\frac{q_{j}^{2}\sigma^{2}}{q_{j}^{2}+\sigma^{2}}.$
With these results in place, we can provide an alternative proof that if $X$ is quadratically convex, then the linear minimax risk for the Gaussian sequence model is equal to linear minimax risk over all rectangular subsets of $X$ (recovering [17, Theorem 7]). Let $\mathcal{R}(X)$ denote the collection of orthosymmetric rectangular subsets of $X$, that is, sets of the form $[-x_{j},x_{j}]_{j\geq 1}\subset X$.
Corollary 3.
Let $X$ be quadratically convex, compact, and orthosymmetric. Then
$\inf_{\widehat{x}=Ay}\sup_{x\in X}\mathbf{E}[\|{\widehat{x}-x}\|_{2}^{2}]=\inf_{A}R^{*}(A,X)=\sup_{H\in\mathcal{R}(X)}\inf_{A}R^{*}(A,H)=\sigma^{2}\sup_{x\in X}\sum_{j\geq 1}\frac{x_{j}^{2}}{x_{j}^{2}+\sigma^{2}}.$
Proof.
Because $X=\mathsf{QHull}(X)$, Corollary 2 implies $\inf_{A}R^{*}(A,X)=\sup_{x\in X}\sum_{j\geq 1}\frac{\sigma^{2}x_{j}^{2}}{x_{j}^{2}+\sigma^{2}}$. For any $x\in X$, the hyperrectangle $H=\bigotimes_{j\geq 1}[-|x_{j}|,|x_{j}|]$ satisfies $\inf_{A}R^{*}(A,H)=\sum_{j\geq 1}\frac{\sigma^{2}x_{j}^{2}}{x_{j}^{2}+\sigma^{2}}$, again by Corollary 2. Take a supremum. ∎
2.1.1 Fundamental limits for the Gaussian sequence model
To introduce the ideas for lower bounds we employ in the remainder of the paper and highlight quadratic convexity, we also review some of the fundamental lower and upper bounds in Gaussian sequence models for general orthosymmetric sets $X$. The following result, which for completeness we prove in Appendix A.4, is typical; it provides worse constants than those available by more careful constructions (cf. [25, Chapter 4]), but it introduces some of our main types of arguments.
Proposition 2.
Let $X$ be an orthosymmetric convex set. Then for any $x\in X$,
$R^{*}(X)\geq\frac{1}{10}\sum_{j\geq 1}x_{j}^{2}\wedge\sigma^{2}.$
As an immediate consequence to Proposition 2, we see that linear estimators are minimax rate optimal whenever $X$ is quadratically convex.
Corollary 4.
Let $X$ be quadratically convex, orthosymmetric, and compact. Then
$R^{*}(X)\leq\inf_{A}R^{*}(A,X)\leq 10R^{*}(X).$
Proof.
Observe that $x_{j}^{2}\wedge\sigma^{2}\geq\frac{x_{j}^{2}\sigma^{2}}{x_{j}^{2}+\sigma^{2}}$, and then apply Corollary 3 and Proposition 2. ∎
2.1.2 Soft-thresholding and nonlinear estimators
Whenever $X$ is quadratically convex, linear estimators are (nearly) minimax optimal, so two fundamental questions remain: first, when do we indeed require nonlinearity, and second, whether there is a generally rate-optimal nonlinear method. Soft-thresholding, which estimates $x$ by elementwise applying the soft-thresholding operator
$\mathsf{S}_{\lambda}(y):=\mathop{\rm sign}(y)\cdot\left({|y|-\lambda}\right)_{+},$
provides this nearly optimal method for Gaussian sequence estimation. Johnstone [25, Corollary 8.4] gives a paradigmatic bound for the risk of soft thresholding on a single coordinate:
Corollary 5.
Let $y\sim\mathsf{N}(x,\sigma^{2})$ and $\delta\in\left({0},{1}\right]$. Then for $\lambda=\sqrt{2\sigma^{2}\log\delta^{-1}}$,
$\mathbf{E}[(\mathsf{S}_{\lambda}(y)-x)^{2}]\leq\delta\sigma^{2}+\left(1+2\log\delta^{-1}\right)\cdot x^{2}\wedge\sigma^{2}.$
If $X\subset\mathbf{R}^{n}$ and $\delta=\frac{1}{n}$, the coordinatewise estimator $\widehat{x}_{j}=\mathsf{S}_{\lambda}(y_{j})$, for $\lambda=\sqrt{2\sigma^{2}\log n}$, obtains
$\mathbf{E}[\|{\widehat{x}-x}\|_{2}^{2}]\leq\sigma^{2}+(2\log n+1)\sum_{j=1}^{n}x_{j}^{2}\wedge\sigma^{2},$
whose risk is essentially optimal (cf. [25, Proposition 8.8] and Proposition 2).
One typically considers the risk in Gaussian sequence models as $\sigma^{2}\to 0$—for example, with the scaling $\sigma^{2}\propto 1/n$ in an $n$-dimensional model—so that when $X$ is “non-pathological” we wish to develop a general estimator that generically achieves optimal scaling, as we will also develop for stochastic optimization in the sequel. For this, we adapt Donoho et al.’s results [17, Thm. 12] by combining projecting coordinates to zero and soft-thresholding. Define
$N(\sigma,X):=\inf\left\{n\in\mathbf{N}~{}\mbox{s.t.}~{}\sup_{x\in X}|x_{j}|\leq\sigma~{}\mbox{for~{}all~{}}j\geq n\right\},$
and for $\lambda=\sqrt{2\sigma^{2}\log N(\sigma,X)}$, define the truncated soft-thresholding estimator
$\widehat{x}_{j}=\begin{cases}\mathsf{S}_{\lambda}(y_{j})&\mbox{if}~{}j\leq N(\sigma,X)\\ 0&\mbox{otherwise}.\end{cases}$ (7)
We have the following corollary, whose proof we defer to appendix A.5.
Corollary 6.
The truncated soft-thresholding estimator (7) satisfies
$\mathbf{E}[\|{\widehat{x}-x}\|_{2}^{2}]\leq\sigma^{2}+(2\log N(\sigma,X)+1)\sum_{j\geq 1}x_{j}^{2}\wedge\sigma^{2}.$
If $N(\sigma,X)$ is polynomial in $1/\sigma$ as $\sigma\to 0$, there exists $C(\sigma)\leq O(1)\log\frac{1}{\sigma}$ such that
$R(\widehat{x},X)\leq C(\sigma)\cdot R^{*}(X).$
The assumption that $N(\sigma,X)$ is polynomial in $1/\sigma$ as $\sigma\to 0$ is typically lenient; any $X\subset\mathbf{R}^{n}$ evidently satisfies it, and so do sets contained in $\ell_{p}$ bodies $\{x\in\mathbf{R}^{\mathbf{N}}\mid\sum_{j=1}^{\infty}a_{j}|x_{j}|^{p}\leq 1\}$ so long as $a_{j}\to\infty$ polynomially quickly in the index $j$. In brief, the (nonlinear) truncated soft-thresholding estimator (7) is nearly minimax rate-optimal: to within a logarithmic factor it achieves the minimax optimal rate for any “sufficiently compact” set $X$. Moreover, because $\frac{1}{2}(a\wedge b)\leq\frac{ab}{a+b}\leq a\wedge b$ for $a,b\geq 0$, the difference between the quantities
$\sup_{x\in X}\sum_{j}x_{j}^{2}\wedge\sigma^{2}~{}~{}\mbox{and}~{}~{}\sup_{x\in\mathsf{QHull}(X)}\sum_{j}x_{j}^{2}\wedge\sigma^{2}$
evidently determines whether nonlinear estimators are necessary.
2.2 Minimax rates for convex stochastic optimization
[…]
$\mathfrak{M}_{k}^{\mathsf{S}}(X,\mathcal{F},\mathcal{P}):=\inf_{\widehat{x}_{k}}\sup_{F\in\mathcal{F}}\sup_{P\in\mathcal{P}}\mathbf{E}\left[f_{P}(\widehat{x}_{k}(S_{1}^{k}))-\inf_{x\in X}f_{P}(x)\right],$
where the expectation is taken over $S_{1}^{k}\stackrel{{\scriptstyle\textup{iid}}}{{\sim}}P$ and the infimum ranges over all measurable functions $\widehat{x}_{k}$ of $\mathcal{S}^{k}$. The average minimax regret provides a related notion, instead taking a supremum over samples $s_{1}^{k}\in\mathcal{S}^{k}$ and measuring instantaneous losses. In this case, an algorithm consists of a sequence of decisions $\widehat{x}_{1},\widehat{x}_{2},\ldots,\widehat{x}_{k}$, where $\widehat{x}_{i}$ is chosen conditional on samples $s_{1}^{i-1}$, so that
$\mathfrak{M}_{k}^{\mathsf{R}}(X,\mathcal{F},\mathcal{S}):=\inf_{\widehat{x}_{1:k}}\sup_{F\in\mathcal{F},s_{1}^{k}\in\mathcal{S}^{k}}\sup_{x\in X}\frac{1}{k}\sum_{i=1}^{k}\left[F\big{(}\widehat{x}_{i}\big{(}s_{1}^{i-1}\big{)},s_{i}\big{)}-F\!\left(x,s_{i}\right)\right].$
In each definition, the point estimates $\widehat{x}$ may lie outside the constraint set $X$, making our lower bounds stronger results; in the language of learning theory [11, 36], we allow “improper” predictions, but compare to the best elements $x\in X$. (In our cases, this does not change regret by more than a constant factor.) As online-to-batch conversions make clear [12], we always have $\mathfrak{M}_{k}^{\mathsf{S}}\leq\mathfrak{M}_{k}^{\mathsf{R}}$; thus we typically provide lower bounds on $\mathfrak{M}_{k}^{\mathsf{S}}$ and upper bounds on $\mathfrak{M}_{k}^{\mathsf{R}}$. In many of the cases we consider, these quantities are essentially equivalent [e.g. 36], though in cases where we wish to provide explicit lower bounds on algorithms we typically use regret.
Lipschitz continuity properties form a central lever for demonstrating convergence in general (potentially non-smooth) stochastic convex optimization [27, 1, 18], and consequently, we study functions for which a norm $\gamma$ on $\mathbf{R}^{n}$ ($\gamma$ as a mnemonic for gradient) specifies these:
$\mathcal{F}^{\gamma,r}:=\left\{F:\mathbf{R}^{n}\times\mathcal{S}\to\mathbf{R}\mid\mbox{for~{}all~{}}x\in\mathbf{R}^{n},~{}g\in\partial_{x}F(x,s),~{}\gamma(g)\leq r\right\},$ (8)
The gradient bound condition $\gamma(g)\leq r$ is equivalent to the Lipschitz condition $|F(x,s)-F(x^{\prime},s)|\leq r\gamma^{*}(x-x^{\prime})$, where $\gamma^{*}$ is the dual norm to $\gamma$. We use the shorthands
[…]
as well as
[…]
as the Lipschitzian properties of $\mathcal{F}$ in relation to $X$ determine the minimax regret and risk. We now discuss the main techniques for lower-bounding the (stochastic) minimax risk, then recapitulate the familiar bounds for online optimization coming from mirror descent and dual averaging, before we move to the more precise results we prove using these techniques and known bounds.
2.2.1 Techniques for minimax optimization lower bounds
We review the Le Cam, Fano and Assouad methods [2, 44, 1, 42] for proving lower bounds. Each reduces estimation to testing then uses information theoretic tools to bound the probability of error in the hypothesis tests.
Le Cam and Fano methods
A standard reduction from estimation to testing (see [18, Ch. 5]) forms the basis for most of our proofs. To set the stage, let $\mathcal{P}$ be a collection of distributions over $\mathcal{S}$ and $L:\mathbf{R}^{n}\times\mathcal{P}\to\mathbf{R}_{+}$ satisfy
$\inf_{x\in X}L(x,P)=0~{}~{}\mbox{for~{}}P\in\mathcal{P}.$
For distributions $P,Q\in\mathcal{P}$, define the separation
$\mathsf{sep}_{L}(P,Q):=\sup\left\{\delta\geq 0\;\middle|\;\;\mbox{for all~{}}x,\begin{array}[]{c}L(x,P)\leq\delta~{}\mbox{implies}~{}L(x,Q)\geq\delta\\ L(x,Q)\leq\delta~{}\mbox{implies}~{}L(x,P)\geq\delta\end{array}\right\}.$
We say a collection $\{P_{v}\}_{v\in\mathcal{V}}\subset\mathcal{P}$ indexed by $\mathcal{V}$ is $\delta$-separated if $\mathsf{sep}_{L}(P_{v},P_{v^{\prime}})\geq\delta$ for $v\neq v^{\prime}$, and any such family induces a testing lower bound:
Lemma 1 (From estimation to testing).
Let $\{P_{v}\}_{v\in\mathcal{V}}\subset\mathcal{P}$ be $\delta$-separated. Then for any joint distribution $\mathbb{P}$ constructed by first drawing $V\in\mathcal{V}$ and then, conditional on $V=v$, sampling $S_{1}^{k}\stackrel{{\scriptstyle\textup{iid}}}{{\sim}}P_{v}$,
$\inf_{\widehat{x}}\sup_{P\in\mathcal{P}}\mathbf{E}_{P}L(\widehat{x}(S_{1}^{k}),P)\geq\delta\inf_{\psi}\mathbb{P}(\psi(S_{1}^{k})\neq V).$
With this, the classical Le Cam and Fano methods are straightforward combinations of Lemma 2 with (respectively) Le Cam’s lemma [44, Lemma 1] and Fano’s inequality [13, Theorem 2.10.1].
Proposition 3 (Le Cam’s method).
Let $P_{0},P_{1}\in\mathcal{P}$ and $\delta>0$ satisfy $\mathsf{sep}_{L}(P_{0},P_{1})\geq\delta$. Then
$\inf_{\widehat{x}}\sup_{P\in\mathcal{P}}\mathbf{E}_{P}L(\widehat{x}(S_{1}^{k}),P)\geq\frac{\delta}{2}\left(1-\big{\|}{P^{k}_{0}-P^{k}_{1}}\big{\|}_{\mathsf{tv}}\right).$
Proposition 4 (Fano’s method).
Let $\mathcal{V}$ be a finite index set for which $\{P_{v}\}_{v\in\mathcal{V}}$ is $\delta$-separated. Then
$\inf_{\widehat{x}}\sup_{P\in\mathcal{P}}\mathbf{E}_{P}L(\widehat{x}(S_{1}^{k}),P)\geq\delta\left(1-\frac{\mathsf{I}(S_{1}^{k};V)+\log 2}{\log|\mathcal{V}|}\right).$
With these tools, minimax lower bounds on the stochastic risk $\mathfrak{M}_{k}^{\mathsf{S}}$ in Section 2 follow by (i) demonstrating an appropriate loss $L$ and (ii) separation. We construct an object that simultaneously accomplishes both of these: for an index set $\mathcal{V}$ with $\{P_{v}\}_{v\in\mathcal{V}}\subset\mathcal{P}$, objective $F$ and population objectives $f_{v}(x):=\mathbf{E}_{P_{v}}[F(x,S)]$, define
$\mathsf{d}_{\textup{opt}}(v,v^{\prime},X):=\inf_{x}\left[f_{v}(x)+f_{v^{\prime}}(x)-\inf_{x\in X}f_{v}(x)-\inf_{x\in X}f_{v^{\prime}}(x)\right].$ (9)
When $\inf_{x\in X}f_{v}(x)=\inf_{x}f_{v}(x)$, which we require because we allow improper predictions, this quantity guarantees appropriate separation. Indeed, under this condition, the loss $L(x,P):=f_{P}(x)-\inf_{x\in X}f_{P}(x)$ satisfies $\inf_{x}L(x,P)=0$ for $P\in\{P_{v}\}_{v\in\mathcal{V}}$. Moreover, if $L(x,P_{v})\leq\frac{1}{2}\mathsf{d}_{\textup{opt}}(v,v^{\prime},X)$, then evidently $L(x,P_{v^{\prime}})\geq\frac{1}{2}\mathsf{d}_{\textup{opt}}(v,v^{\prime},X)$, and so $\mathsf{sep}_{L}(P_{v},P_{v^{\prime}})\geq\frac{1}{2}\mathsf{d}_{\textup{opt}}(v,v^{\prime},X)$. Lemma 1 then implies the following result, which is essentially present in the paper [1] (cf. [18]).
Lemma 2 (From optimization to function estimation).
Let $\{P_{v}\}_{v\in\mathcal{V}}\subset\mathcal{P}$ be a collection of distributions on $\mathcal{S}$ and $\mathcal{F}$ be a collection a functions $\mathbf{R}^{n}\times\mathcal{S}\to\mathbf{R}$, where for some $F\in\mathcal{F}$, the optimization distance (9) satisfies $\mathsf{d}_{\textup{opt}}(v,v^{\prime},X)\geq\delta\geq 0$ for all $v\neq v^{\prime}\in\mathcal{V}$. Then
$\mathfrak{M}_{k}^{\mathsf{R}}(X,\mathcal{F})\geq\mathfrak{M}_{k}^{\mathsf{S}}(X,\mathcal{F},\mathcal{P})\geq\mathfrak{M}_{k}^{\mathsf{S}}(X,\{F\},\mathcal{P})\geq\frac{\delta}{2}\inf_{\psi}\mathbb{P}(\psi(S_{1}^{k})\neq V).$
Our general strategy for proving lower bounds on $\mathfrak{M}_{k}^{\mathsf{S}}$ is as follows:
  • Choose a function $F\in\mathcal{F}$ and define $\mathcal{V}$ and $\{P_{v}\}_{v\in\mathcal{V}}\subset\mathcal{P}$ such that $\mathsf{d}_{\textup{opt}}(v,v^{\prime},X)\geq\delta>0$.
  • Lower bound the testing error $\inf_{\psi}\mathbb{P}(\psi(S_{1}^{k})\neq V)$, and choose the largest separation $\delta$ to make this testing error a positive constant.
To showcase this proof technique, we prove that minimax stochastic risk for $1$-dimensional optimization has lower bound $1/\sqrt{k}$; we use this to address technicalities in later proofs.
Lemma 3.
[…]
$\mathfrak{M}_{k}^{\mathsf{S}}([-1,1],\mathcal{F}^{n=1})\geq\frac{1}{4\sqrt{6k}}.$
Proof.
Let $X=[-1,1]$ and $\mathcal{S}=\{\pm 1\},\mathcal{V}=\{\pm 1\}$. To see the separation condition, let $F(x,s):=|x-s|$. For $\delta\in[0,\frac{1}{2}]$, we define $P_{v}$ so that for $S\sim P_{v}$,
$P_{v}(S=v)=\frac{1+\delta}{2},~{}~{}P_{v}(S=-v)=\frac{1-\delta}{2}.$
Then $f_{v}(x)=\frac{1+\delta}{2}|x-v|+\frac{1-\delta}{2}|x+v|$ and $\inf_{x}f_{v}(x)=\inf_{x\in[-1,1]}f_{v}(x)=\frac{1-\delta}{2}$. To lower bound the separation, note that
$f_{1}(x)+f_{-1}(x)-\inf_{X}f_{1}-\inf_{X}f_{-1}=|x-1|+|x+1|-(1-\delta)\geq\delta.$
This yields $\mathsf{d}_{\textup{opt}}(1,-1,X)\geq\delta$.
We lower bound the testing error via Proposition 3: by Pinsker’s inequality that $\left\|{P-Q}\right\|_{\mathsf{tv}}^{2}\leq\frac{1}{2}D_{\textup{kl}}\left({P}|\!|{Q}\right)$ for any distributions $P$ and $Q$, and that $D_{\textup{kl}}\left({P^{k}}|\!|{Q^{k}}\right)=kD_{\textup{kl}}\left({P}|\!|{Q}\right)$, we obtain
$\inf_{\psi:\mathcal{S}^{k}\to\{\pm 1\}}\mathbb{P}(\psi(S_{1}^{k})\neq V)=\frac{1}{2}\left(1-\big{\|}{P_{1}^{k}-P_{-1}^{k}}\big{\|}_{\mathsf{tv}}\right)\geq\frac{1}{2}\left(1-\sqrt{\frac{k}{2}D_{\textup{kl}}\left({P_{1}}|\!|{P_{-1}}\right)}\right).$
Noting that $D_{\textup{kl}}\left({P_{1}}|\!|{P_{-1}}\right)=\delta\log\frac{1+\delta}{1-\delta}\leq 3\delta^{2}$ for $\delta\in[0,\frac{1}{2}]$, setting $\delta=1/\sqrt{6k}$ yields the result. ∎
The Assouad method
Assouad’s method [2] reduces the problem of estimation (or optimization) to one of multiple binary hypothesis tests. In this case, we index a set of distributions $\mathcal{P}=\{P_{v}\}_{v\in\mathcal{V}}$ on a set $\mathcal{S}$ by the hypercube $\mathcal{V}=\{\pm 1\}^{n}$. For a function $F:\mathbf{R}^{n}\times\mathcal{S}\to\mathbf{R}$, we define $f_{v}(x):=\mathbf{E}_{P_{v}}[F(x,S)]$. Then for a vector $\delta\in\mathbf{R}^{n}_{+}$, following Duchi [18, Lemma 5.3.2], we say that the functions $\{f_{v}\}$ induce a $\delta$-separation in Hamming metric if
$f_{v}(x)-\inf_{x\in X}f_{v}(x)\geq\sum_{j=1}^{n}\delta_{j}1\!\left({\mathop{\rm sign}(x_{j})\neq v_{j}}\right).$ (10)
Lemma 4 (Generalized Assouad’s method).
Let $S_{1}^{k}\stackrel{{\scriptstyle\textup{iid}}}{{\sim}}P_{V}$, where $V\sim\mathsf{Uniform}(\{\pm 1\}^{n})$. Define the averages
$\mathbb{P}_{+j}:=\frac{1}{2^{n-1}}\sum_{v:v_{j}=1}P_{v}^{k}~{}~{}\mbox{and}~{}~{}\mathbb{P}_{-j}:=\frac{1}{2^{n-1}}\sum_{v:v_{j}=-1}P_{v}^{k}.$
Assume that the collection $\{f_{v}\}$ for $f_{v}=\mathbf{E}_{P_{v}}[F(\cdot,S)]$ induces a $\delta$-separation (10). Then
$\mathfrak{M}_{k}^{\mathsf{S}}(X,\{F\},\mathcal{P})\geq\frac{1}{2}\sum_{j=1}^{n}\delta_{j}\left(1-\left\|{\mathbb{P}_{+j}-\mathbb{P}_{-j}}\right\|_{\mathsf{tv}}\right).$
2.2.2 Stochastic gradient methods, mirror descent, and regret
Let us briefly review the canonical algorithms for solving the problem (1) and their associated convergence guarantees. For an algorithm outputing points $x_{1},\ldots,x_{k}$, the regret on the sequence $F(\cdot,s_{i})$ with respect to a point $x$ is
$\mathsf{Regret}_{k}(x):=\sum_{i=1}^{k}[F(x_{i},s_{i})-F(x,s_{i})].$
Recalling the definition $\mathrm{D}_{h}(x,x_{0})=h(x)-h(x_{0})-\left\langle{\nabla h(x_{0})},{x-x_{0}}\right\rangle$ of the Bregman divergence, the mirror descent algorithm [27, 5] iteratively sets
$g_{i}\in\partial_{x}F(x_{i},s_{i})~{}~{}\mbox{and~{}updates}~{}~{}x^{\mathsf{MD}}_{i+1}:=\mathop{\textup{argmin}}_{x}\left\{g_{i}^{\top}x+\frac{1}{\alpha}\mathrm{D}_{h}(x,x_{i})\right\}$ (11)
[…]
$\mathsf{Regret}_{k}(x)\leq\frac{\mathrm{D}_{h}(x,x_{0})}{\alpha}+\frac{\alpha}{2}\sum_{i\leq k}\left\|{g_{i}}\right\|_{*}^{2}~{}~{}\mbox{for all}~{}x\in\mathop{\rm dom}h.$ (12)
The choice $h(x)=\frac{1}{2}\left\|{x}\right\|_{2}^{2}$ recovers the classical stochastic gradient method, while the $p$-norm algorithms [23, 34, 28, 18], defined for $1<p\leq 2$, use $h(x)=\frac{1}{2(p-1)}\left\|{x}\right\|_{p}^{2}$; each is strongly convex with respect to the $\ell_{p}$-norm $\left\|{\cdot}\right\|_{p}$. If $G=\{g\in\partial_{x}F(x,s)\mid x\in X,s\in\mathcal{S}\}$ denotes the set of possible subgradients, the regret guarantee (12) becomes
$\mathsf{Regret}_{k}(x)\leq\frac{\left\|{x}\right\|_{p}^{2}}{2(p-1)\alpha}+\frac{k\alpha}{2}\sup_{g\in G}\left\|{g}\right\|^{2}_{q}$
if $x_{0}=0$ and $q=\frac{p}{p-1}$ is conjugate to $p$. The choice $\alpha=\frac{1}{\sqrt{k}}\sup_{x\in X}\left\|{x}\right\|_{p}/\sup_{\gamma(g)\leq 1}\left\|{g}\right\|_{q}$ gives the following now standard minimax regret bound [cf. 35, Corollary 2.18].
Proposition 5.
Let $X$ be closed convex orthosymmetric, $\gamma$ a norm, and $1<p\leq 2$, $q=\frac{p}{p-1}$. Mirror descent with $h(x):=\frac{1}{2(q-1)}\left\|{x}\right\|_{p}^{2}$ and stepsize $\alpha=\frac{\sup_{x\in X}\left\|{x}\right\|_{p}}{\sqrt{k(p-1)}\sup_{\gamma(g)\leq 1}\left\|{g}\right\|_{q}}$ achieves regret
$\mathfrak{M}_{k}^{\mathsf{R}}(X,\gamma)\leq\frac{\sup_{x\in X}\left\|{x}\right\|_{p}\sup_{g\in\mathbf{B}_{\gamma}(0,1)}\|g\|_{q}}{\sqrt{k(p-1)}}.$
As we state in our definitions of minimax risk and regret, we allow the point estimates to lie outside the constraint set $X$, though as with the dual averaging update (5), one may incorporate constraints by replacing $h$ with $h+\mathbf{1}_{X}$. Even without such constraints, the form (12) and Proposition 5 still capture the regret for all common constraint sets $X$ [34]. Example 1 ($p$-norm algorithms):  Let $n\geq 3$. With $h(x)=\frac{1}{2(p-1)}\left\|{x}\right\|_{p}^{2}$, taking $p=1+\frac{1}{\log n}$ and $q=1+\log n$, we have $\left\|{g}\right\|_{q}\leq e\left\|{g}\right\|_{\infty}$ for all $g\in\mathbf{R}^{n}$, while $\left\|{x}\right\|_{p}\leq\left\|{x}\right\|_{1}$. Assuming that gradients $g$ belong to the $\ell_{\infty}$-ball of radius 1, inequality (12) and Proposition 5 show that for any $X\subset\mathbf{R}^{n}$, the stepsize $\alpha=e^{-1}\sqrt{\log n/k}\sup_{x\in X}\left\|{x}\right\|_{1}$ yields
$\sup_{x\in X}\mathsf{Regret}_{k}(x)\leq e\sqrt{k\log n}\sup_{x\in X}\left\|{x}\right\|_{1}.$
Taking the comparator class $X$ to be the $\ell_{1}$-ball gives regret $e\cdot\sqrt{k\log n}$. $\Diamond$
2.2.3 Euclidean gradient methods and quadratic convexity
We frequently focus on distance generating functions of the form $h(x)=\frac{1}{2}\left\langle{x},{Ax}\right\rangle$ for a fixed positive semi-definite matrix $A$. For an arbitrary $A$, we will refer to these methods as Euclidean gradient methods and for a diagonal $A$ as diagonally scaled gradient methods. In this case, the mirror descent update is the stochastic gradient update with ${A}^{-1}g$, where $g$ is a stochastic subgradient. We refer to all such methods as methods of linear type, as their update sequence guarantees the linearity
$x_{k}=-A^{-1}\sum_{i=1}^{k-1}g_{i}.$
With updates of this type, convex quadratic hulls of both the domain $X$ and the collections of gradients $\partial F(x,s)$ appear in convergence bounds (the sequel shows this appearance is no accident).
The diagonally scaled gradient methods (componentwise re-scaling of the subgradients) are equivalent to using $h_{\Lambda}(x):=\frac{1}{2}x^{\top}\Lambda x$ for $\Lambda=\mathrm{diag}(\lambda)\succeq 0$ in the mirror descent update (11). In this case, for any norm $\gamma$ on the gradients, the minimax regret bound (12) becomes
$\sup_{x\in X}\mathsf{Regret}_{k,\Lambda}(x)\leq\frac{1}{2k}\left[\sup_{x\in X}x^{\top}\Lambda x+\sum_{i\leq k}g_{i}^{\top}\Lambda^{-1}g_{i}\right]\leq\frac{1}{2k}\left[\sup_{x\in X}x^{\top}\Lambda x+k\sup_{g\in\mathbf{B}_{\gamma}(0,1)}g^{\top}\Lambda^{-1}g\right].$
The rightmost term upper bounds the minimax regret, so we may take an infimum over $\Lambda$, yielding
$\mathfrak{M}_{k}^{\mathsf{R}}(X,\gamma)\leq\frac{1}{2k}\inf_{\lambda\succeq 0}\sup_{x\in X}\sup_{g\in\mathbf{B}_{\gamma}(0,1)}\bigg{[}\sum_{j\leq n}\lambda_{j}x_{j}^{2}+k\sum_{j\leq n}\frac{1}{\lambda_{j}}g^{2}_{j}\bigg{]}$ (13)
The regret bound (13) holds without assumptions on $X$ or $\gamma$. The key is that strong duality allows us to relate this quantity to the quadratic convex hulls (recall Sec. 2) of $X$ and $\mathbf{B}_{\gamma}(0,1)$, which will in turn allow us to prove minimax optimality of Euclidean procedures in the coming sections.
Proposition 6.
Let $G,X\subset\mathbf{R}^{n}$ be compact convex sets. Then
$\inf_{\lambda\succ 0}\sup_{x\in X,g\in G}\left\{\lambda^{\top}x^{2}+\left(\frac{1}{\lambda}\right)^{\top}g^{2}\right\}=\sup_{x\in\mathsf{QHull}(X),g\in\mathsf{QHull}(G)}\inf_{\lambda\succ 0}\left\{\lambda^{\top}x^{2}+\left(\frac{1}{\lambda}\right)^{\top}g^{2}\right\}.$
Proof.
The (weighted) squared $2$-norm becomes a linear functional when lifted to the squared sets $X^{2}:=\{x^{2}\mid x\in X\}$ and $G^{2}$. Indeed, defining $J(\tau,w,\lambda):=\lambda^{\top}\tau+(\frac{1}{\lambda})^{\top}w$, the function $J$ is concave-convex: it is linear (a fortiori concave) in $(\tau,w)$ and convex in $\lambda\succ 0$. Thus, using that the set $\{\lambda\in\mathbf{R}^{n}_{+}\}$ is convex and $\mathsf{SqHull}(X)\times\mathsf{SqHull}(G)$ is convex compact (because $X$ and $G$ compact), Sion’s minimax theorem [37] implies
$\inf_{\lambda\succ 0}\sup_{x\in X,g\in G}\left\{\lambda^{\top}x^{2}+\left(\frac{1}{\lambda}\right)^{\top}g^{2}\right\}$ $=\inf_{\lambda\succ 0}\sup_{\tau\in\mathsf{SqHull}(X),w\in\mathsf{SqHull}(G)}\left\{\lambda^{\top}\tau+\left(\frac{1}{\lambda}\right)^{\top}w\right\}$
$=\sup_{\tau\in\mathsf{SqHull}(X),w\in\mathsf{SqHull}(G)}\inf_{\lambda\succ 0}\left\{\lambda^{\top}\tau+\left(\frac{1}{\lambda}\right)^{\top}w\right\}.$
Replacing $\tau$ with $x^{2}$ and $w$ with $v^{2}$ gives the result. ∎
Proposition 6 provides a powerful hammer for diagonally scaled Euclidean algorithms, as we can choose an optimal scaling for any fixed pair $x,g$, taking a worst case over such pairs:
Corollary 7.
Let $X$ be a convex compact set and $\gamma$ a norm. Then
$\mathfrak{M}_{k}^{\mathsf{R}}(X,\gamma)\leq\frac{1}{\sqrt{k}}\sup_{g\in\mathsf{QHull}(\mathbf{B}_{\gamma}(0,1))}\sup_{x\in\mathsf{QHull}(X)}x^{\top}g,$
and diagonally scaled gradient methods achieve this regret.
Proof.
Apply Proposition 6 to the regret (13) and use $\inf_{\lambda>0}a\lambda+b/\lambda=2\sqrt{ab}$ for $a,b\geq 0$. ∎
Corollary 7 allows concrete upper bounds on minimax risk and regret, and it is suggestive that the quadratic hulls of (respectively) the comparator class $X$ and subgradient sets relate essentially to convergence rates, similarly to the case with Gaussian sequence models. The remainder of the paper develops the analogy of linear and nonlinear updates in stochastic optimization problems with those in Gaussian sequence models, highlighting when methods of linear type are minimax rate optimal, and when more computational power—we require nonlinearity—is necessary.
3 Minimax optimality and quadratically convex constraint sets
We begin by providing lower bounds on the minimax risk and matching upper bounds on the minimax regret of convex optimization over quadratically convex constraint sets, where diagonally scaled gradient methods achieve the regret bounds. While the analogy with the Gaussian sequence model is nearly complete, in distinction to the work of Donoho et al. (where results depend solely on the constraints $X$, as in Corollary 3), our results necessarily depend on the geometry of the subdifferential. Consequently, we distinguish throughout this section between quadratically and non-quadratically convex geometry of the gradients. To set the stage for our contributions, we begin with the classical case of $X=\mathbf{B}_{p}(0,1)$ with $p\in[2,\infty]$ (so that $X$ is quadratically convex) and norm $\gamma=\|{\cdot}\|_{r}$ with $r\geq 1$. We then turn to arbitrary quadratically convex constraint sets and first show results in the case of general quadratically convex norms on the subgradients. We conclude the section by proving that, when the subgradients do not lie in a quadratically convex set but lie in a weighted $\ell_{r}$ ball (for $r\in[1,2]$), diagonally scaled gradient methods are still minimax rate optimal.
3.1 A warm-up: $p$-norm constraint sets for $p\geq 2$
Though the results for the basic case that $X$ is an $\ell_{p}$-ball while the gradients belong to a different $\ell_{r}$-ball are special cases of the theorems to come, the proofs provide intuition for the later results. We distinguish between two cases depending on the value of $r$ in the gradient norm. The case that $r\in[1,2]$ corresponds roughly to “sparse” gradients, while the case $r\geq 2$ corresponds to harder problems with dense gradients. We provide information theoretic proofs of the following two results in Sec. 3.1.1 and 3.1.2, respectively.
Proposition 7 (Sparse gradients).
Let $X=\mathbf{B}_{p}(0,1)$ with $p\geq 2$ and $\gamma(\cdot)=\left\|{\cdot}\right\|_{r}$ where $r\in[1,2]$. Then
$1\wedge\frac{n^{\frac{1}{2}-\frac{1}{p}}}{\sqrt{k}}\lesssim\mathfrak{M}_{k}^{\mathsf{S}}(X,\gamma)\leq\mathfrak{M}_{k}^{\mathsf{R}}(X,\gamma)\lesssim 1\wedge\frac{n^{\frac{1}{2}-\frac{1}{p}}}{\sqrt{k}}.$
Proposition 8 (Dense gradients).
Let $X=\mathbf{B}_{p}(0,1)$ with $p\geq 2$ and $\gamma(\cdot)=\|\cdot\|_{r}$ with $r\geq 2$. Then
$1\wedge\frac{n^{\frac{1}{2}-\frac{1}{p}}n^{\frac{1}{2}-\frac{1}{r}}}{\sqrt{k}}\lesssim\mathfrak{M}_{k}^{\mathsf{S}}(X,\gamma)\leq\mathfrak{M}_{k}^{\mathsf{R}}(X,\gamma)\lesssim 1\wedge\frac{n^{\frac{1}{2}-\frac{1}{p}}n^{\frac{1}{2}-\frac{1}{r}}}{\sqrt{k}}.$
In both cases, the stochastic gradient method achieves the regret upper bound via a straightforward application of the regret bounds (12) with $h(x)=\frac{1}{2}\left\|{x}\right\|_{2}^{2}$ or Corollary 7; a method of linear type is optimal.
3.1.1 Proof of Proposition 7
We use the general information-theoretic framework of reduction from estimation to testing presented in Section 2.2.1 to prove the lower bound. Define the constant $c_{p}=n^{-1/p}$, which satisfies $c_{p}\left\|{\mathbf{1}}\right\|_{p}=1$ as $\left\|{\mathbf{1}}\right\|_{p}=n^{1/p}$.
Separation
Consider the sample space $\mathcal{S}=\{\pm e_{j}\}_{j\leq n}$ and the function
$F(x,s):=\langle|s|,|x-c_{p}s|\rangle=\sum_{j=1}^{n}|s_{j}||x_{j}-c_{p}s_{j}|$
so $F$ belongs to $\mathcal{F}^{\gamma,1}$ as $\partial_{x}F(x,s)\subset\cup_{j=1}^{n}[-1,1]\cdot e_{j}$ when $s\in\mathcal{S}$. Letting $\delta\in[0,1/2]$ to be determined, for $v\in\{\pm 1\}^{n}$, we define $P_{v}$ such that for $S\sim P_{v}$ we have
$S=\begin{cases}v_{j}e_{j}&~{}~{}\mbox{with probability}~{}~{}\frac{1+\delta}{2n}\\ -v_{j}e_{j}&~{}~{}\mbox{with probability}~{}~{}\frac{1-\delta}{2n}.\end{cases}$
We then have
$f_{v}(x)=\frac{1}{2n}\sum_{j=1}^{n}(1+\delta v_{j})|x_{j}-c_{p}|+(1-\delta v_{j})|x_{j}+c_{p}|,$
so $\mathop{\textup{argmin}}_{x}f_{v}(x)=c_{p}v\in\mathbf{B}_{p}(0,1)$, whence
$f_{v}^{*}:=\inf f_{v}=\inf_{x\in X}f_{v}(x)=(1-\delta)c_{p}.$
where $p^{\ast}$ is such that $11/p+1/p^{\ast}=$ 1. For $v,v^{\prime}\in\{\pm 1\}^{n}$, we have
$f_{v}(x)+f_{v^{\prime}}(x)$
$=\frac{1}{n}\sum_{j:v_{j}=v_{j}^{\prime}}\left[(1+\delta v_{j})|x_{j}-c_{p}|+(1-\delta v_{j})|x_{j}+c_{p}|\right]+\frac{1}{n}\sum_{j:v_{j}\neq v^{\prime}_{j}}\left[|x_{j}-c_{p}|+|x_{j}+c_{p}|\right],$
whose minimizer is $c_{p}(v+v^{\prime})/2\in\mathbf{B}_{p}(0,1)$. The optimization distance then satisfies
$\mathsf{d}_{\textup{opt}}(v,v^{\prime},X)=\frac{2c_{p}}{n}[(n-\mathsf{d}_{\textup{Ham}}(v,v^{\prime}))(1-\delta)+\mathsf{d}_{\textup{Ham}}(v,v^{\prime})]-2(1-\delta)c_{p}=2c_{p}\delta\frac{\mathsf{d}_{\textup{Ham}}(v,v^{\prime})}{n},$
where $\mathsf{d}_{\textup{Ham}}(v,v^{\prime})$ is the Hamming distance between $v$ and $v^{\prime}$. The Gilbert-Varshimov bound [42, Example 5.3] guarantees the existence of a packing set $\mathcal{V}\subset\{\pm 1\}^{n}$ of cardinality at least $\exp(n/8)$ satisfying $\mathsf{d}_{\textup{Ham}}(v,v^{\prime})\geq\frac{n}{4}$ for each $v\neq v^{\prime}\in\mathcal{V}$. Then
$\mathsf{d}_{\textup{opt}}(v,v^{\prime},X)\geq\frac{c_{p}\delta}{2}=\frac{1}{2}n^{-1/p}\delta~{}~{}\mbox{for~{}all~{}}v\neq v^{\prime}\in\mathcal{V},$
and applying Lemma 2 yields
$\mathfrak{M}_{k}^{\mathsf{S}}(X,\gamma)\geq\frac{\delta}{4}n^{-1/p}\inf_{\psi}\mathbb{P}(\psi(S_{1}^{k})\neq V).$
Bounding the testing error
We bound the testing error with Fano’s inequality and via the trivial upper bound $\mathsf{I}(S_{1}^{k};V)\leq\max_{v,v^{\prime}}D_{\textup{kl}}\left({P_{v}^{k}}|\!|{P_{v^{\prime}}^{k}}\right)$ on the mutual information [42, Ch. 14]. Using the identity $\delta\log\frac{1+\delta}{1-\delta}\leq 3\delta^{2}$ for $\delta\leq\frac{1}{2}$, we have $D_{\textup{kl}}\left({P_{v}}|\!|{P_{v^{\prime}}}\right)\leq 3\delta^{2}$, and so Fano’s inequality (Proposition 4) and our choice that $|\mathcal{V}|\geq\exp(n/8)$ imply
$\inf_{\psi}\mathbb{P}(\psi(S_{1}^{k})\neq V)\geq\left(1-\frac{3k\delta^{2}+\log 2}{n/8}\right).$
In the case that $n\geq 32\log 2$, choosing $\delta=\sqrt{\frac{n}{48k}}$ yields the desired lower bound, once we recognize that $\mathfrak{M}_{k}^{\mathsf{S}}$ is non-increasing in $k$. In the case that $n<32\log 2$, with $\mathcal{F}^{n=1}$ as in Lemma 3, that any $1$-dimensional optimization problem may be embedded into a $n$-dimensional problem yields
$\mathfrak{M}_{k}^{\mathsf{S}}(X,\gamma)\geq\mathfrak{M}_{k}^{\mathsf{S}}([-1,1],\mathcal{F}^{n=1})\gtrsim\frac{1}{\sqrt{k}}.$
This gives the lower bound for all $n\in\mathbf{N}$.
To conclude the proof, we establish an upper bound on the minimax regret. Because $p\geq 2$, $\mathsf{QHull}(X)=X=\mathbf{B}_{p}(0,1)$, while for $1\leq r\leq 2$, $\left\|{g}\right\|_{r}\leq\left\|{g}\right\|_{2}$ and so $\overline{G}:=\mathsf{QHull}(\mathbf{B}_{\gamma}(0,1))\subset\mathbf{B}_{2}(0,1)$. Then $\sup_{x\in X}\sup_{g\in\overline{G}}x^{\top}g\leq\sup_{x\in X}\left\|{x}\right\|_{2}=n^{\frac{1}{2}-\frac{1}{p}}$.
3.1.2 Proof of Proposition 8
The proof is similar to Proposition 7, so we forego some of the details, again letting $c_{p}=n^{-1/p}$ so that $c_{p}\mathbf{1}\in\mathbf{B}_{p}(0,1)$.
Separation
Let $\eta>0$ to be determined, and consider the sample space $\mathcal{S}=\{\pm 1\}^{n}$ and objectives
$F(x,s):=\eta\left\|{x-c_{p}s}\right\|_{1}.$
For $v\in\{\pm 1\}^{n}$, we define $P_{v}$ so that $S\sim P_{v}$ has independent coordinates satisfying
$P_{v}(S_{j}=v_{j})=\frac{1+\delta}{2},~{}~{}P_{v}(S_{j}=-v_{j})=\frac{1-\delta}{2}.$
This yields
$f_{v}(x)=\frac{\eta(1+\delta)}{2}\left\|{x-c_{p}v}\right\|_{1}+\frac{\eta(1-\delta)}{2}\left\|{x+c_{p}v}\right\|_{1},$
and $\mathop{\textup{argmin}}_{x}f_{v}(x)=c_{p}v\in\mathbf{B}_{p}(0,1)$, with $f_{v}^{*}=nc_{p}\eta(1-\delta)$. For $v\neq v^{\prime}$, a quick calculation yields
$(f_{v}+f_{v^{\prime}})^{*}=\inf_{x}(f_{v}(x)+f_{v^{\prime}}(x))=2\eta c_{p}\mathsf{d}_{\textup{Ham}}(v,v^{\prime})+2\eta c_{p}(1-\delta)(n-\mathsf{d}_{\textup{Ham}}(v,v^{\prime})).$
Then considering again the Gilbert-Varshimov packing $\mathcal{V}\subset\{\pm 1\}^{n}$, we have
$\mathsf{d}_{\textup{opt}}(v,v^{\prime},X)=2\eta c_{p}\delta\mathsf{d}_{\textup{Ham}}(v,v^{\prime})\geq\frac{1}{2}\eta\cdot n^{1-1/p}~{}~{}\mbox{for~{}}v\neq v^{\prime}.$
Bounding the testing error
Note that for $\delta\leq\frac{1}{2}$,
$D_{\textup{kl}}\left({P_{v}}|\!|{P_{v^{\prime}}}\right)=\sum_{j:v_{j}\neq v_{j}^{\prime}}\delta\log\frac{1+\delta}{1-\delta}\leq 3n\delta^{2},$
so $\mathsf{I}(S_{1}^{k};V)\leq 3kn\delta^{2}$. For $F$ to remain in $\mathcal{F}^{\gamma,1}$, we must have that $\eta\left\|{\mathbf{1}}\right\|_{r}\leq 1$; noting that $\left\|{\mathbf{1}}\right\|_{r}=n^{1/r}$, we choose $\eta=n^{-1/r}$. In the case that $n\geq 32\log 2$, taking $\delta=1/\sqrt{48k}$ yields the minimax lower-bound
$\mathfrak{M}_{k}^{\mathsf{S}}(X,\gamma)\gtrsim\frac{n^{1-\frac{1}{p}}n^{-\frac{1}{r}}}{\sqrt{k}}=\frac{n^{\frac{1}{2}-\frac{1}{p}}n^{\frac{1}{2}-\frac{1}{r}}}{\sqrt{k}}.$
In the case that $n<32\log 2$, we once again refer Lemma 3, which concludes the proof for the lower bound on the minimax stochastic risk.
The upper bound again follows immediately via Corollary 7, as both $X$ and $\mathbf{B}_{\gamma}(0,1)$ are quadratically convex, and for $q=\frac{p}{p-1}$,
$\sup_{\left\|{x}\right\|_{p}\leq 1}\sup_{\left\|{g}\right\|_{r}\leq 1}x^{\top}g=\sup_{\left\|{g}\right\|_{r}\leq 1}\left\|{g}\right\|_{q}=n^{\frac{1}{q}-\frac{1}{r}}=n^{1-\frac{1}{p}-\frac{1}{r}}.$
3.2 General quadratically convex constraints
We now turn to the more general case that $X$ is an arbitrary orthosymmetric quadratically convex body. Our techniques build out of the ideas of Donoho et al. [17] in Gaussian sequence estimation, where, as in Section 2.1 the largest hyperrectangle in $X$ governs the performance of linear estimators. As in the previous section, we divide our analysis into cases, depending on the quadratic convexity of the gradient norm $\gamma$ (analogs of $r\lessgtr 2$ in Propositions 7 and 8). Our starting point is a general lower bound relying on rectangular structures in the primal $X$ and dual gradient spaces. For the proposition, we use a specialization of the function families (8) to rectangular sets, where for $M\in\mathbf{R}_{+}^{n}$ we define
$\mathcal{F}^{M}:=\Big{\{}F:\mathbf{R}^{n}\times\mathcal{S}\to\mathbf{R}\mid\mbox{for~{}all~{}}x\in\mathbf{R}^{n}~{}\mbox{and}~{}g\in\partial_{x}F(x,s),~{}\max_{j\leq n}\frac{|g_{j}|}{M_{j}}\leq 1\Big{\}}.$
Then Duchi et al. [20] provide the following lower bound.
Proposition 9 ([20], Proposition 1).
Let $M\in\mathbf{R}_{+}^{n}$ and $\mathcal{F}^{M}$ be as above. Let $a\in\mathbf{R}_{+}^{n}$ and assume the hyperrectangular containment $\prod_{j=1}^{n}[-a_{j},a_{j}]\subset X$. Then
$\mathfrak{M}_{k}^{\mathsf{S}}(X,\mathcal{F}^{M})\geq\frac{1}{8\sqrt{k\log 3}}\sum_{j=1}^{n}M_{j}a_{j}.$
An immediate extension of this result provides lower bounds on the minimax stochastic risk for all orthosymmetric norms $\gamma$, by which we mean $\gamma(\sigma\odot v)=\gamma(v)$ for all sign vectors $\sigma$, and convex sets $X$ without any restriction on qudaratic convexity. (It will be the matching upper bounds that require this.)
Corollary 8.
Let $X\subset\mathbf{R}^{n}$ be an orthosymmetric convex set and $\gamma$ an orthosymmetric norm. Then
$\mathfrak{M}_{k}^{\mathsf{S}}(X,\gamma)\geq\frac{1}{8\sqrt{k\log 3}}\sup_{x\in X}\gamma^{*}(x).$
Proof.
Define the hyperrectangle $\mathsf{Rec}(x):=\prod_{j\leq n}[-|x_{j}|,|x_{j}|]$, so that $X\supset\mathsf{Rec}(x)$ for all $x\in X$ by orthosymmetry. Additionally, recalling the notation (8) of $\mathcal{F}^{\gamma,1}$ and $\mathcal{F}^{M}$, if $M\in\mathbf{R}^{n}_{+}$ satisfies $\gamma(M)\leq 1$ then, by orthosymmetry of $\gamma$, $\mathcal{F}^{\gamma,1}\supset\mathcal{F}^{M}$. Thus
$\mathfrak{M}_{k}^{\mathsf{S}}(X,\gamma)\geq\mathfrak{M}_{k}^{\mathsf{S}}(\mathsf{Rec}(x),\gamma)\geq\mathfrak{M}_{k}^{\mathsf{S}}(\mathsf{Rec}(x),\mathcal{F}^{M})\geq\frac{1}{8\sqrt{k\log 3}}\sum_{j\leq n}|x_{j}|M_{j}$
for all $M\in\mathbf{B}_{\gamma}(0,1)\cap\mathbf{R}_{+}^{n}$ and $x\in X$. Taking a supremum over $M\in\mathbf{B}_{\gamma}(0,1)$ and $x\in X$, we have
$\mathfrak{M}_{k}^{\mathsf{S}}(X,\gamma)\geq\frac{1}{8\sqrt{k\log 3}}\sup_{x\in X}\sup_{\gamma(M)\leq 1}x^{\top}M=\frac{1}{8\sqrt{k\log 3}}\sup_{x\in X}\gamma^{*}(x)$
as desired. ∎
3.2.1 Orthosymmetric and quadratically convex gradient norms
We now provide lower bounds on minimax risk complementary to Corollary 7, focusing first on the case that the gradient norm $\gamma$ is quadratically convex.
Assumption A1.
The norm $\gamma$ is orthosymmetric and quadratically convex, meaning $\gamma(\sigma\odot v)=\gamma(v)$ for all $\sigma\in\{\pm 1\}^{n}$ and $\mathbf{B}_{\gamma}(0,1)$ is quadratically convex.
With this, we have the following theorem, which shows that diagonally-scaled gradient methods are minimax rate optimal, and that the constants are sharp up to a factor of $9$, whenever the gradient norms are quadratically convex. While the constant $9$ is looser than that Donoho et al. [17] provide for Gaussian sequence models, this theorem highlights the essential structural similarity between the sequence model case and stochastic optimization methods.
Theorem 1.
Let Assumption A1 hold and let $X$ be quadratically convex, orthosymmetric, and compact. Then
$\frac{1}{8\sqrt{\log 3}}\frac{1}{\sqrt{k}}\sup_{x\in X}\gamma^{*}(x)\leq\mathfrak{M}_{k}^{\mathsf{S}}(X,\gamma)\leq\mathfrak{M}_{k}^{\mathsf{R}}(X,\gamma)\leq\frac{1}{\sqrt{k}}\sup_{x\in X}\gamma^{*}(x).$
There exists $\lambda^{\ast}\in\mathbf{R}^{n}_{+}$ such that diagonally scaled gradient methods with $\lambda^{\ast}$ achieve this rate.
Proof.
For the upper bound, we use Corollary 7. Because $\mathbf{B}_{\gamma}(0,1)$ is quadratically convex, we have $\mathsf{QHull}(\mathbf{B}_{\gamma}(0,1))=\mathbf{B}_{\gamma}(0,1)$, so that $\sup_{g\in\mathsf{QHull}(\mathbf{B}_{\gamma}(0,1))}x^{\top}g=\gamma^{*}(x)$. Then note that $\mathsf{QHull}(X)=X$ as well. The lower bound follows immediately from Corollary 8. ∎
3.2.2 Arbitrary gradient norms
When the norm $\gamma$ on the gradients defines a non-quadratically convex norm ball $\mathbf{B}_{\gamma}(0,1)$—for example, when the gradients belong to an $\ell_{r}$-norm ball for $r\in[1,2]$—our results become less general. Nonetheless, when $\gamma$ is a weighted $\ell_{r}$-norm ball (for $r\in[1,2]$), diagonally scaled gradient methods remain minimax rate optimal, as Corollary 9 will show. When the norms $\gamma$ are arbitrary we use the rescaled vector $\textup{res}(x,\gamma):=(x_{j}/\gamma(e_{j}))_{j=1}^{n}$, where $e_{j}$ are the standard basis vectors:
Theorem 2.
Let $X$ be an orthosymmetric, quadratically convex, convex and compact set and $\gamma$ an arbitrary norm. For any $d\in\mathbf{N}$,
$\frac{1}{8\sqrt{k\log 3}}\left(1-\frac{d}{k\log 3}\right)\sup_{x\in X,\left\|{x}\right\|_{0}\leq d}\left\|{\textup{res}(x,\gamma)}\right\|_{2}$ $\leq\mathfrak{M}_{k}^{\mathsf{S}}(X,\gamma)$
$\leq\mathfrak{M}_{k}^{\mathsf{R}}(X,\gamma)\leq\frac{1}{\sqrt{k}}\sup_{x\in X}\sup_{g\in\mathsf{QHull}(\mathbf{B}_{\gamma}(0,1))}x^{\top}g.$
Corollary 7 gives the upper bound in the theorem. The lower bound consists of an application of Assouad’s method (Lemma 4), but, in parallel to the warm-up examples (Propositions 7 and 8), we construct well-separated functions with “sparse” gradients. See Appendix B.1 for a proof.
We can develop a corollary of this result when the norm $\gamma$ is a weighted-$\ell_{r}$ norm for $r\in[1,2]$. While these do not induce quadratically convex norm balls, the previous theorem still guarantees that diagonally scaled gradient methods are minimax rate optimal.
Corollary 9.
Let the conditions of Theorem 2 hold and assume that $\gamma(g)=\left\|{\beta\odot g}\right\|_{r}$ with $r\in[1,2]$, $\beta_{j}>0$ and $(\beta\odot g)_{j}=\beta_{j}g_{j}$. Then for $k\geq 2n$,
$\frac{1}{16}\frac{1}{\sqrt{k}}\sup_{x\in X}\left\|{\textup{res}(x,\gamma)}\right\|_{2}\leq\mathfrak{M}_{k}^{\mathsf{S}}(X,\gamma)\leq\mathfrak{M}_{k}^{\mathsf{R}}(X,\gamma)\leq\frac{1}{\sqrt{k}}\sup_{x\in X}\left\|{\textup{res}(x,\gamma)}\right\|_{2}.$
There exists $\lambda^{\ast}\in\mathbf{R}^{n}_{+}$ such that diagonally scaled gradient methods with $\lambda^{\ast}$ achieve this rate.
A minor modification of Theorem 2 gives the lower bound, while we obtain the upper bound by noting that the quadratic hull of a weighted-$\ell_{r}$ norm ball for $r\in[1,2]$ is the weighted-$\ell_{2}$ norm ball. The dual norm of $\gamma(g)=\left\|{\beta\odot g}\right\|_{2}$ being $\gamma^{\ast}(g)=\left\|{g/\beta}\right\|_{2}$, the upper bound holds by duality. See Appendix B.2 for the (short) proof.
Theorem 1 and Corollary 9 show that for a large collection of norms $\gamma$ on the gradients, diagonally scaled gradient methods areminimax rate optimal. Arguing that diagonally scaled gradient methods are minimax rate optimal when $\gamma$ is neither a weighted-$\ell_{r}$ norm nor induces a quadratically convex unit ball remains an open question, though weighted-$\ell_{r}$ norms for $r\in[1,\infty]$ cover the majority of practical applications of stochastic gradient methods.
4 Beyond quadratic convexity: the necessity of non-linear methods
For $X\subset\mathbf{R}^{n}$ quadratically convex, the results in Section 3 show that methods of linear type achieve optimal rates of convergence. When the constraint set is not quadratically convex, it is unclear whether methods of linear type are sufficient to achieve optimal rates. As we now show, they are not: we exhibit collections of problem instances in which the constraint sets are orthosymmetric convex bodies but not quadratically convex, and where methods of linear type must have regret at least a factor $\sqrt{n/\log n}$ worse than the minimax optimal rate, which (non-linear) mirror descent with appropriate distance generating function achieves. We also develop more general results to highlight the way in which the quadratic hull of the underlying constraint set $X$ necessarily characterizes the regret of Euclidean gradient methods, which allows for a more explicit delineation of those sets $X$ for which nonlinear methods are necessary: when $\sup_{x\in X}\left\|{x}\right\|_{1}$ is much smaller than $\sup_{x\in\mathsf{QHull}(X)}\left\|{x}\right\|_{1}$.
To construct these problem instances, we first turn to simple non-quadratically convex constraint sets: $\ell_{p}$ balls for $p\in[1,2]$. We measure subgradient norms in the dual $\ell_{p^{*}}$ norm, $p^{*}=\frac{p}{p-1}$. Our analysis consists of two steps: we first prove sharp minimax rates on these problem instances and show that mirror descent with the right (non-linear) distance generating function is minimax rate optimal. These results extend those of Agarwal et al. [1], who provide matching lower and upper bounds for $p\geq 1+c$ for a fixed numerical constant $c>0$. In contrast, we prove sharp minimax rates for all $p\geq 1$. To precisely characterize the gap between linear and non-linear methods, we show that for any linear pre-conditioner, we can exhibit functions for which the regret of Euclidean gradient methods is nearly the simple upper regret bound of standard gradient methods, Eq. (12) with $h(x)=\frac{1}{2}\left\|{x}\right\|_{2}^{2}$. Thus, when $p$ is very close to $2$ (nearly quadratically convex), the gap remains within a constant factor, whereas when $p$ is close to $1$, the gap can be as large as $\sqrt{n/\log n}$.
4.1 Minimax rates for $p$-norm constraint sets and general convex bodies
For $p\in[1,2]$, we consider the constraint set $X=\mathbf{B}_{p}(0,1)\subset\mathbf{R}^{n}$ and bound gradients with norm $\gamma=\|\cdot\|_{p^{\ast}}$. We begin by proving sharp minimax rates on this collection of problems and show that, in these cases, non-linear mirror descent is minimax optimal.
Theorem 3.
Let $p\in[1,2]$, $X=\mathbf{B}_{p}(0,1)\subset\mathbf{R}^{n}$ and $\gamma=\|\cdot\|_{p^{\ast}}$.
  1. (i)
    If $1\leq p\leq 1+1/\log(2n)$, then
    $1\wedge\sqrt{\frac{\log(2n)}{k}}\lesssim\mathfrak{M}_{k}^{\mathsf{S}}(X,\gamma)\leq\mathfrak{M}_{k}^{\mathsf{R}}(X,\gamma)\lesssim 1\wedge\sqrt{\frac{\log(2n)}{k}}.$
  2. (ii)
    If $1+1/\log(2n)<p\leq 2$, then
    $1\wedge\sqrt{\frac{1}{k(p-1)}}\lesssim\mathfrak{M}_{k}^{\mathsf{S}}(X,\gamma)\leq\mathfrak{M}_{k}^{\mathsf{R}}(X,\gamma)\lesssim 1\wedge\sqrt{\frac{1}{k(p-1)}}$
In both cases, mirror descent (11) with distance generating function $h(x):=\frac{1}{2(a-1)}\left\|{x}\right\|_{a}^{2}$ for $a=\max\{1+\frac{1}{\log(2n)},p\}$ achieves the optimal rate.
The upper bound essentially follows from Proposition 5, and the lower bound uses reductions from estimation to testing. See Appendix C.1 for a proof.
We also provide a few results extending to potentially infinite-dimensional spaces. To reduce complexity, we focus on the case that the gradients are bounded in $\ell_{\infty}$-norm—the “most” quadratically convex set—to make dependence on $X$ the clearest. In the sequel it will sometimes be useful to consider sequence space $\mathbf{R}^{\mathbf{N}}$, so we give a result that allows the infinite dimensional containment $X\subset\mathbf{R}^{\mathbf{N}}$; in this case we consider domains $X$ that are appropriately compact for the $\ell_{1}$-norm. A sufficient condition for the results is that $X$ have finite effective dimension, which we define by
$\textup{effdim}(X):=\inf_{\beta\geq 0}\left\{e^{1/\beta}\mid\sup_{x\in X}\sum_{j=1}^{\infty}j^{\beta}|x_{j}|\leq e\sup_{x\in X}\left\|{x}\right\|_{1}\right\},$ (14)
where $\textup{effdim}(X)=+\infty$ if only $\beta=0$ satisfies the inequality. It is immediate that if $X\subset\mathbf{R}^{n}$, then $\textup{effdim}(X)\leq n$ because $n^{1/\log n}=e$. The next observation summarizes a few sufficient conditions for the effective dimension (14) to be finite. We include the proof in Appendix C.2 for completeness.
Observation 4.1.
The following conditions are sufficient to guarantee $\textup{effdim}(X)<\infty$.
  1. i.
    If $X\subset\mathbf{R}^{n}$, then $\textup{effdim}(X)\leq n$.
  2. ii.
    If $\lim_{\beta\downarrow 0}\sup_{x\in X}\sum_{j\geq 1}j^{\beta}|x_{j}|<\infty$, then $\textup{effdim}(X)<\infty$.
  3. iii.
    Let $N<\infty$ and $\gamma>0$ satisfy the tail condition that $\sup_{x\in X}\sum_{j=1}^{N}|x_{j}|\geq n^{2\gamma}\sup_{x\in X}\sum_{j>n}|x_{j}|$ for all $n>N$. Then $\textup{effdim}(X)\leq O(1)\cdot(N^{2}+\exp(\frac{3}{\gamma}\log\frac{1}{\gamma}))$.
The next proposition then shows that the $\ell_{1}$-diameter of $X$ always provides lower bounds on the (stochastic) minimax risk, while $p$-norm-based mirror descent algorithms achieve regret at most logarithmic in the effective-dimension (14), making the $\ell_{1}$-diameter of $X$ the central quantity governing minimax risk.
Proposition 10.
Let $X$ be an orthosymmetric convex set. Then
$\frac{1}{8\sqrt{\log 3}}\frac{1}{\sqrt{k}}\sup_{x\in X}\left\|{x}\right\|_{1}\leq\mathfrak{M}_{k}^{\mathsf{S}}(X,\left\|{\cdot}\right\|_{\infty}).$
If additionally $X$ has finite effective dimension $N=\textup{effdim}(X)$ as in (14), then
$\mathfrak{M}_{k}^{\mathsf{R}}(X,\left\|{\cdot}\right\|_{\infty})\leq O(1)\cdot\sqrt{\log N}\frac{1}{\sqrt{k}}\sup_{x\in X}\left\|{x}\right\|_{1}.$
Letting $\beta=\frac{1}{\log N}$ and defining the operator $A$ by $Ax=(j^{\beta}x_{j})_{j\geq 1}$, the mirror descent method with distance generating function $h(x)=\frac{1}{2(p-1)}\left\|{Ax}\right\|_{p}^{2}$ for $p=\frac{2}{2-\beta}$ achieves this regret.
To preview our discussion to come, note the similarities between Proposition 10 and Corollary 6 in the Gaussian sequence model case: so long as $X$ is appropriately regular, there is a “universal” nonlinear method—in the case of Gaussian sequence models, soft-thresholding with truncation (7), in the case of stochastic and online optimization, mirror descent with a $p$-norm-based distance-generating function—that achieves nearly rate-optimal minimax risk. The analogy extends spiritually as well, as in both Gaussian sequence models and stochastic optimization [25, 16, 17, 27, 5], the universal methods were originally designed for high-dimensional signals $x$.
Proof.
The lower bound is a direct consequence of Corollary 8. To demonstrate the upper bound, we take $X\subset\mathbf{R}^{\mathbf{N}}$, as the special case of finite dimensions follows immediately. For notational simplicity, for a positive sequence $(a_{j})_{j\in\mathbf{N}}$ to be chosen we let $A:\mathbf{R}^{\mathbf{N}}\to\mathbf{R}^{\mathbf{N}}$ be the diagonal linear operator $Ax=(a_{j}x_{j})_{j\in\mathbf{N}}$. For $p\in\left({1},{2}\right]$ to be chosen as well, define the distance generating function
$h(x)=\frac{1}{2(p-1)}\left\|{Ax}\right\|_{p}^{2}.$
By an analogous argument to the finite-dimensional case, this is strongly convex with respect to the norm $\left\|{x}\right\|=\left\|{Ax}\right\|_{p}$ and has dual norm $\left\|{g}\right\|_{*}=\left\|{A^{-1}g}\right\|_{q}$ for $q=\frac{p}{p-1}$. For any sequence of subgradients $g_{i}\in[-1,1]^{\mathbf{N}}$, we thus obtain regret bound
$\mathsf{Regret}_{k}(x)\leq\frac{1}{2(p-1)\alpha}\left\|{Ax}\right\|_{p}^{2}+\frac{\alpha}{2}\sum_{i=1}^{k}\left\|{A^{-1}g_{i}}\right\|_{q}^{2}.$
Notably, for $g\in[-1,1]^{\mathbf{N}}$ we have
$\left\|{A^{-1}g}\right\|_{q}^{q}\leq\sum_{j=1}^{\infty}a_{j}^{-q}.$
We use the assumption (14) that $X$ has finite effective dimension $N$, so that for $\beta=\frac{1}{\log N}\leq 1$, $\sup_{x\in X}\sum_{j=1}^{\infty}j^{\beta}|x_{j}|\leq e\sup_{x\in X}\left\|{x}\right\|_{1}$. Letting $a_{j}=j^{\beta}$, we take $q=\frac{2}{\beta}$, whence $\|{A^{-1}g}\|_{q}^{q}=\sum_{j=1}^{\infty}j^{-2}=\frac{\pi^{2}}{6}$. As this also gives conjugate $p=\frac{q}{q-1}=1+\frac{\beta}{2-\beta}$, we obtain the regret bound
$\mathsf{Regret}_{k}(x)\leq\frac{1}{\alpha}\frac{2-\beta}{2\beta}\left\|{Ax}\right\|_{p}^{2}+O(1)\cdot\alpha k\leq O(1)\left[\frac{1}{\alpha\beta}\sup_{x\in X}\left\|{x}\right\|_{1}^{2}+\alpha k\right].$
Take $\alpha=\sup_{x\in X}\left\|{x}\right\|_{1}/\sqrt{k\beta}$ to obtain the convergence upper bound. ∎
We make one final remark on these results, and include some commentary in the final discussion on future work. Corollary 8 shows that
$\frac{1}{8\sqrt{k\log 3}}\sup_{x\in X}\sup_{g:\gamma(g)\leq 1}x^{\top}g\leq\mathfrak{M}_{k}^{\mathsf{S}}(X,\gamma)$
for arbitrary orthosymmetric convex $X$ and norms $\gamma$. This is sharp whenever $\gamma$ is an $\ell_{q}$ norm for $2\leq q\leq\infty$, as then mirror descent with distance generating function $h(x)=\frac{1}{2(p-1)}\left\|{x}\right\|_{p}^{2}$, $p=\max\{\frac{q}{q-1},1+\frac{1}{\log n}\}$ achieves the minimax lower bound, with average regret scaling at worst as $\sqrt{\log n}\sup_{x\in X}\left\|{x}\right\|_{q}/\sqrt{k}$. It is also not a tight lower bound in general: when $X=[-1,1]^{n}$ and $\gamma(g)=\left\|{g}\right\|_{1}$, then Proposition 7 shows that $\sqrt{n/k}$ is the correct scaling—a factor $\sqrt{n}$ larger than $\sup_{x,g}x^{\top}g$.
4.2 Hard problems for Euclidean gradient methods and quadratic hulls
Theorem 3 shows that (non-linear) mirror descent methods are minimax rate-optimal for $\ell_{p}$-ball constraint sets, $p\in[1,2]$, with gradients contained in the corresponding dual $\ell_{q}$-norm ball ($q=\frac{p}{p-1}$). For such problems and $p$, standard subgradient methods achieve worst-case regret $O(n^{1/2-1/q}/\sqrt{k})$. This upper bound is in fact sharp: in the next theorem, we show that for any method of linear type, we can construct a sequence of (linear) functions such that the method’s regret achieves the worst-case upper bound for standard subgradient methods, precisely quantifying the gap between linear and non-linear methods for this problem class. To that end, let
$\mathsf{Regret}_{k,A}(x):=\sum_{i=1}^{k}g_{i}^{\top}(x_{i}-x)$
denote the regret of the Euclidean online mirror descent method with distance generating function $h_{A}(x)=\frac{1}{2}x^{\top}Ax$ for functions $F_{i}$ with subgradients $g_{i}$. In the lower bounds to come, we take $F_{i}(x)=g_{i}^{\top}x$ to be linear, so that $\nabla F_{i}(x)=g_{i}$ is independent of $x$.
Theorem 4.
For any $A\succeq 0$ and $p\in[1,2]$ with $q=\frac{p}{p-1}$, there exists a sequence of vectors $g_{i}\in\mathbf{R}^{n}$, $\left\|{g_{i}}\right\|_{q}\leq 1$, and point $x\in\mathbf{R}^{n}$ with $\left\|{x}\right\|_{p}\leq 1$ such that
$\mathsf{Regret}_{k,A}(x)\geq\frac{1}{2}\min\left\{k/2,\sqrt{2k}\cdot n^{1/2-1/q}\right\}.$
Scaled identity matrices $A=c\cdot I_{n}$ achieve these bounds to within a factor of $\sqrt{2}$ for $k\geq 2n^{1-2/q}$.
Proof.
Let $A\succ 0$ be a positive semi-definite matrix for the distance generating function $h_{A}(x)=\frac{1}{2}x^{\top}Ax$ defined above, and let $q=\frac{p}{p-1}$ be the conjugate to $p$. We choose linear functions $F_{i}(x):=g_{i}^{\top}x$ where $g_{i}\in\mathbf{B}_{q}(0,1)$. In this case, letting $\{x_{i}\}_{i\leq k}$ be the points mirror descent plays, the regret with respect to $x\in\mathbf{R}^{n}$ is
$\mathsf{Regret}_{k,A}(x)=\sum_{i\leq k}F_{i}(x_{i})-F_{i}(x)=\sum_{i\leq k}g_{i}^{\top}(x_{i}-x),$
so that
$\mathsf{Regret}^{\ast}_{k,A}:=\sup_{\left\|{x}\right\|_{p}\leq 1}\mathsf{Regret}_{k,A}(x)=\bigg{\|}{\sum_{i\leq k}g_{i}}\bigg{\|}_{q}+\frac{1}{2}\sum_{i\leq k}\left\|{g_{i}}\right\|_{A^{-1}}^{2}-\frac{1}{2}\bigg{\|}{\sum_{i\leq k}g_{i}}\bigg{\|}_{A^{-1}}^{2}.$
Now, we choose linear functions $f_{i}$ so that the regret is large. To do so, choose vectors
$u\in\mathop{\textup{argmax}}_{\left\|{s}\right\|_{q}\leq 1}s^{\top}A^{-1}s~{}~{}\mbox{and}~{}~{}v\in\mathop{\textup{argmin}}_{\left\|{s}\right\|_{q}=1}s^{\top}A^{-1}s.$ (15)
Then set the (gradient) vectors $g_{i}\in\mathbf{R}^{n}$ so that for a $\delta\in[0,1]$ to be chosen,
  1. (a)
    $g_{i}=u$ for $k/4$ of the indices $i\in[k]$
  2. (b)
    $g_{i}=-u$ for $k/4$ of the indices $i\in[k]$
  3. (c)
    $g_{i}=v$ for $\frac{k}{4}(1+\delta)$ of the indices $i\in[k]$
  4. (d)
    $g_{i}=-v$ for $\frac{k}{4}(1-\delta)$ of the indices $i\in[k]$.
With these choices, we obtain the regret lower bound
$\mathsf{Regret}^{\ast}_{k,A}$ $\geq\sup_{\delta\leq 1}\left[\frac{k}{2}\delta\left\|{v}\right\|_{q}+\frac{k}{4}u^{\top}{A}^{-1}u-\frac{\delta^{2}k^{2}}{8}v^{\top}{A}^{-1}v\right]$
$\geq\frac{k}{4}\cdot\left[u^{\top}A^{-1}u+\min\left\{1,\frac{2\left\|{v}\right\|_{q}}{kv^{\top}A^{-1}v}\right\}\left\|{v}\right\|_{q}\right].$ (16)
We now consider two cases. In the first, $A$ is large enough that $\left\|{v}\right\|_{q}\geq\frac{1}{2}kv^{\top}A^{-1}v$. Then the regret bound (16) becomes
$\mathsf{Regret}^{*}_{k,A}\geq\frac{k}{4}\left[u^{\top}A^{-1}u+\left\|{v}\right\|_{q}\right]\geq\frac{k}{4},$
as $\left\|{v}\right\|_{q}=1$ by the construction (15). This gives the first result of the theorem. For the second claim, which holds in the case that $\left\|{v}\right\|_{q}<\frac{1}{2}kv^{\top}A^{-1}v$, we consider the operator norms of general invertible linear operators. For a mapping $T:\mathbf{R}^{n}\to\mathbf{R}^{n}$, define the $\ell_{p}$ to $\ell_{q}$ operator norm
$\left\|{T}\right\|_{\ell_{p}\to\ell_{q}}:=\sup_{x\neq 0}\frac{\left\|{T(x)}\right\|_{q}}{\left\|{x}\right\|_{p}}.$
Then the construction (15) evidently yields
$u^{\top}A^{-1}u=\|{A^{-1/2}}\|_{\ell_{q}\to\ell_{2}}^{2}~{}~{}~{}\mbox{and}~{}~{}~{}\frac{\left\|{v}\right\|_{q}^{2}}{v^{\top}A^{-1}v}=\sup_{x\neq 0}\frac{\|{A^{1/2}x}\|_{q}^{2}}{\left\|{x}\right\|_{2}^{2}}=\|{A^{1/2}}\|_{\ell_{2}\to\ell_{q}}^{2}.$
Revisiting the regret (16), we obtain
$\mathsf{Regret}^{*}_{k,A}\geq\frac{k}{4}\cdot\left[\left\|{A^{-1/2}}\right\|_{\ell_{q}\to\ell_{2}}^{2}+\frac{2}{k}\left\|{A^{1/2}}\right\|_{\ell_{2}\to\ell_{q}}^{2}\right]\geq\sqrt{\frac{k}{2}}\|{A^{-1/2}}\|_{\ell_{q}\to\ell_{2}}\|{A^{1/2}}\|_{\ell_{2}\to\ell_{q}},$
where we have used that $ab\leq\frac{1}{2}a^{2}+\frac{1}{2}b^{2}$ for all $a,b$. But for any invertible linear operator, standard results on the Banach-Mazur distance [41, Corollary 2.3.2] imply that
$\inf_{A\succ 0}\left\|{A}\right\|_{\ell_{2}\to\ell_{q}}\left\|{A^{-1}}\right\|_{\ell_{q}\to\ell_{2}}\geq n^{1/2-1/q}.$
This gives the lower bound.
For the claimed upper bound, note for $h(x)=\frac{1}{2\alpha}\left\|{x}\right\|_{2}^{2}$ (i.e. $A=\frac{1}{\alpha}I_{n}$) and initial point $x_{0}=\mathbf{0}$, we have $\mathsf{Regret}_{k}(x)\leq\frac{1}{2\alpha}\left\|{x}\right\|_{2}^{2}+\frac{\alpha}{2}\sum_{i=1}^{k}\left\|{g_{i}}\right\|_{2}^{2}$. As $\left\|{x}\right\|_{2}\leq 1$ whenever $\left\|{x}\right\|_{p}\leq 1$ and and $\left\|{g}\right\|_{2}\leq n^{1/2-1/q}$ whenever $\left\|{g}\right\|_{q}\leq 1$, we have $\mathsf{Regret}_{k}(x)\leq\frac{1}{2\alpha}+\frac{\alpha}{2}kn^{1-2/q}$. Choose $\alpha=(kn^{1-2/q})^{-1/2}$ to minimize this bound and achieve $\sup_{\left\|{x}\right\|_{p}\leq 1}\mathsf{Regret}_{k}(x)\leq\sqrt{k}n^{1/2-1/q}$. ∎
These results explicitly exhibit a gap between methods of linear type and non-linear mirror descent methods for this problem class. In contrast to the frequent practice in literature of simply comparing regret upper bounds—prima facie illogical—we demonstrate the gap indeed must hold.
In combination with Theorem 4, Proposition 5 precisely separates linear and non-linear mirror descent on these problems for all values of $p\in[1,2]$. Indeed, when $p=1$, for any pre-conditioner $A$, there exists a problem on which Euclidean gradient methods have regret at least $\Omega(1)\sqrt{n/k}$. On the same problem, non-linear mirror descent has regret at most $O(1)\sqrt{\log n/k}$, showing the advertised $\sqrt{n/\log n}$ gap. When $p\geq 2-1/\log n$ (so $X$ is nearly quadratically convex), the gap reduces to at most a constant factor.
The quadratically convex hulls of both the domain $X$ and the space $G:=\{\partial F(x;s)\}_{x\in X,s\in\mathcal{S}}$ of potential subgradients turn out to completely govern the behavior of linear-type methods. In this sense, Corollary 7 is sharp, even to the leading constant factor: the cost (in a minimax sense) of using linear methods is that regret necessarily scales with $\mathsf{QHull}(X)$ and $\mathsf{QHull}(G)$.
Theorem 5.
Let $X\subset\mathbf{R}^{n}$ and $G\subset\mathbf{R}^{n}$ be orthosymmetric convex bodies. For any sequence $A(k)\succeq 0$, there exist sequences of vectors $g_{i}=g_{i}(k)\in G$, $i=1,\ldots,k$, such that
$\liminf_{k}\frac{1}{\sqrt{k}}\sup_{x\in X}\mathsf{Regret}_{k,A(k)}(x)\geq\sup_{x\in\mathsf{QHull}(X)}\sup_{g\in\mathsf{QHull}(G)}x^{\top}g.$
Additionally, for each $k$ there exists a diagonal matrix $D$ such that for any sequence of convex functions $F_{i}:X\to\mathbf{R}$ for which $\partial F_{i}\subset G$,
$\frac{1}{\sqrt{k}}\sup_{x\in X}\mathsf{Regret}_{k,D}(x)\leq\sup_{x\in\mathsf{QHull}(X)}\sup_{g\in\mathsf{QHull}(G)}x^{\top}g.$
See Appendix D for the proof of the result, which parallels that of Theorem 4.
To make a neater analogy with Gaussian sequence models $y=x+\xi$ for $\xi\sim\mathsf{N}(0,\sigma^{2}I_{n})$, where the scale of noise $\xi_{j}$ on different coordinates is identical, we specialize to the case that $\gamma(g)=\left\|{g}\right\|_{\infty}$, that is, $G=\mathbf{B}_{\infty}(0,1)$. Recalling the notation (8) identifying $\mathcal{F}^{\left\|{\cdot}\right\|_{\infty},1}$ as those functions whose subgradients $g\in\partial F(x,s)$ belong to the $\ell_{\infty}$-ball, the next corollary then follows immediately.
Corollary 10.
Let $X\subset\mathbf{R}^{n}$ be an orthosymmetric convex body. For any sequence $A(k)\succeq 0$, there exist sequences of vectors $g_{i}=g_{i}(k)\in\mathbf{R}^{n}$, $i=1,\ldots,k$, with $\left\|{g_{i}}\right\|_{\infty}\leq 1$, such that
$\liminf_{k}\frac{1}{\sqrt{k}}\sup_{x\in X}\mathsf{Regret}_{k,A(k)}(x)\geq\sup_{x\in\mathsf{QHull}(X)}\left\|{x}\right\|_{1}.$
Additionally, for each $k$ there exists a diagonal matrix $D$ such that for any sequence of convex functions $F_{i}\in\mathcal{F}^{\left\|{\cdot}\right\|_{\infty},1}$,
$\sup_{x\in X}\mathsf{Regret}_{k,D}(x)\leq\sqrt{k}\sup_{x\in\mathsf{QHull}(X)}\left\|{x}\right\|_{1}.$
In brief, the regret of Euclidean gradient methods necessarily scales with the size of the quadratic hull $\mathsf{QHull}(X)$. Contrasting this result with Proposition 10, we see with nonlinear methods, the regret need scale only as $\sup_{x\in X}\left\|{x}\right\|_{1}$ rather than as $\sup_{x\in\mathsf{QHull}(X)}\left\|{x}\right\|_{1}$, so that the gap between convergence achievable by linear and nonlinear methods is large precisely when
$\frac{\sup_{x\in\mathsf{QHull}(X)}\left\|{x}\right\|_{1}}{\sup_{x\in X}\left\|{x}\right\|_{1}}\gg 1.$
5 The role of quadratic convexity in sequence models and first-order methods
The results in Section 2.1 highlight and recapitulate some of the ways that quadratic convexity distinguishes linear and nonlinear methods in Gaussian sequence models. Theorems 4 and 5, along with the complimentary results in Proposition 10, address these differences for stochastic optimization problems. So in both sequence models and convex optimization, geometric aspects of the underlying set $X$ determine nonlinear methods’ necessity. The analogy between sequence models and stochastic optimization methods is not perfect, however, as there are sets $X$ for which linear methods are minimax rate optimal for stochastic optimization problems and not for sequence models and vice versa. In both problem families, a particular “distance” of a set $X$ from quadratic convexity delineates determines when nonlinear methods are necessary; we show that these can be different.
We begin by translating the results in Section 2.1 on Gaussian sequence models into a more geometric form; Donoho et al. [17] more or less give this translation but we make a few minor modifications for convenience in exposition. The measure of size most natural for Gaussian sequence models turns out to be (duality-based variants of) the Kolmogorov $n$-width of the underlying set:
Definition 5.1.
The $n$-width of a set $X$ is
$0pt^{2}(n):=\inf_{\mathbf{0}\preceq d\preceq\mathbf{1},\langle\mathbf{1},d\rangle=n}\sup_{x\in X}\sum_{j}(1-d_{j})x_{j}^{2}.$
The nonlinear $n$-width of $X$ is
$w_{\textup{nl}}^{2}(n):=\sup_{x\in X}\inf_{\mathbf{0}\preceq d\preceq\mathbf{1},\langle\mathbf{1},d\rangle=n}\sum_{j}(1-d_{j})x_{j}^{2}.$
Recalling Corollaries 2 and 6, the gap between the linear minimax and nonlinear minimax risk is large for (compact) convex sets $X$ whenever the difference between
$\sup_{x\in X}\sum_{j=1}^{\infty}\frac{x_{j}^{2}\sigma^{2}}{x_{j}^{2}+\sigma^{2}}~{}~{}\mbox{and}~{}~{}\sup_{q\in\mathsf{QHull}(X)}\sum_{j=1}^{\infty}\frac{q_{j}^{2}\sigma^{2}}{q_{j}^{2}+\sigma^{2}}.$
The next proposition, parts of which are present in Donoho et al. [17, Section 6], connects the $n$-widths to the linear and nonlinear minimax risks, where for a vector $x\in\mathbf{R}^{\mathbf{N}}$, we let $|x_{(1)}|\geq|x_{(2)}|\geq\cdots$ denote the entries of $x$ sorted by magnitude.
Proposition 11.
For any compact orthosymmetric convex set $X$,
$0pt^{2}(n)=\sup_{q\in\mathsf{QHull}(X)}\sum_{j\geq n+1}q_{(j)}^{2}~{}~{}\mbox{and}~{}~{}w_{\textup{nl}}^{2}(n)=\sup_{x\in X}\sum_{j\geq n+1}x_{(j)}^{2}.$
Additionally, for any $\sigma^{2}$,
$\sup_{x\in X}\sum_{j}x_{j}^{2}\wedge\sigma^{2}=\inf_{n}\left\{w_{\textup{nl}}^{2}(n)+n\sigma^{2}\right\},$
and
$\sup_{q\in\mathsf{QHull}(X)}\sum_{j}q_{j}^{2}\wedge\sigma^{2}=\inf_{n}\left\{0pt^{2}(n)+n\sigma^{2}\right\}.$
Proof.
For the characterizations, we use duality to see that
$0pt^{2}(n)$ $=\inf_{\mathbf{0}\preceq d\preceq\mathbf{1},\langle\mathbf{1},d\rangle=n}\sup_{v\in\mathsf{SqHull}(X)}\langle\mathbf{1}-d,v\rangle$
$=\sup_{v\in\mathsf{SqHull}(X)}\inf_{\mathbf{0}\preceq d\preceq\mathbf{1},\langle\mathbf{1},d\rangle=n}\langle\mathbf{1}-d,v\rangle=\sup_{v\in\mathsf{SqHull}(X)}\sum_{j\geq n+1}v_{(j)},$
while it is immediate that $w_{\textup{nl}}^{2}(n)=\sup_{x\in X}\sum_{j\geq n+1}x_{(j)}^{2}$. Then we recognize that for any sorted nonnegative vector $a_{1}\geq a_{2}\geq\cdots$,
$\inf_{n}\bigg{\{}\sum_{j\geq n+1}a_{j}+n\sigma^{2}\bigg{\}}=\inf_{n}\bigg{\{}\sum_{j\geq n+1}a_{j}+\sum_{j=1}^{n}\sigma^{2}\bigg{\}}=\sum_{j}a_{j}\wedge\sigma^{2}$
by choosing any $n$ such that $a_{j}\leq\sigma^{2}$ for all $j\geq n$, while $a_{j}\geq\sigma^{2}$ for $j<n$. ∎
Using Proposition 11, we see that under the conditions of Corollary 6, because for any $a,b\geq 0$ we have $\frac{1}{2}\min\{a,b\}\leq\frac{ab}{a+b}\leq\min\{a,b\}$, the linear sequence model risk satisfies
$\frac{1}{2}\inf_{n}\left\{0pt^{2}(n)+n\sigma^{2}\right\}\leq R_{\textup{lin}}^{*}(X)\leq\inf_{n}\left\{0pt^{2}(n)+n\sigma^{2}\right\},$
while by Proposition 2, the (nonlinear) minimax risk satisfies
$\frac{1}{10}\inf_{n}\left\{w_{\textup{nl}}^{2}(n)+n\sigma^{2}\right\}\leq R^{*}(X)\lesssim\log\frac{1}{\sigma^{2}}\cdot\inf_{n}\left\{w_{\textup{nl}}^{2}(n)+n\sigma^{2}\right\}.$
The linear and nonlinear $n$-widths of $X$ therefore (up to a logarithmic factor in $\frac{1}{\sigma}$) determine the risk in sequence models, so that when they are similar the linear and nonlinear minimax risks coincide. In stochastic optimization, in contrast, Section 4 shows that convergence guarantees for methods of linear type coincide with those for arbitrary methods (again, up to logarithmic factors) if and only if
$\sup_{x\in\mathsf{QHull}(X)}\left\|{x}\right\|_{1}\asymp\sup_{x\in X}\left\|{x}\right\|_{1}.$
The typical scenario in sequence models one considers is the risk as $\sigma\downarrow 0$, and the following essentially trivial observation shows that for regular enough sets $X$, when the linear and nonlinear $n$-widths differ, the rates at which $R_{\textup{lin}}^{*}$ and $R^{*}$ converge to zero differ.
Observation 5.1.
Let $\alpha>0$ be such that $n^{\alpha}w_{\textup{nl}}^{2}(n)\asymp 0pt^{*}>0$ as $n\to\infty$, and assume for some $\beta>0$ that $0pt^{2}(n)\geq n^{\beta}w_{\textup{nl}}^{2}(n)$. Then
$\frac{R^{*}(X)}{R_{\textup{lin}}^{*}(X)}\lesssim\log\frac{1}{\sigma^{2}}\cdot\sigma^{\frac{2\beta}{(1+\alpha)(1+\alpha-\beta)}}\to 0~{}~{}\mbox{as}~{}~{}\sigma\downarrow 0.$
Proof.
The assumption that $w_{\textup{nl}}^{2}(n)\lesssim 0pt^{*}n^{-\alpha}$ guarantees the assumptions of Corollary 6 apply, and we necessarily have $\beta\leq\alpha$ (as otherwise we would have $w_{\textup{nl}}^{2}(n)\to\infty$). Thus there exists a (numerical) constant $C<\infty$ such that for all small enough $\sigma^{2}>0$,
$R^{*}(X)\leq C\log\frac{1}{\sigma^{2}}\cdot\inf_{n}\left\{0pt^{*}n^{-\alpha}+n\sigma^{2}\right\}\asymp\log\frac{1}{\sigma^{2}}\cdot\sigma^{\frac{2\alpha}{1+\alpha}},$
which follows by setting $n=\sigma^{-2/(1+\alpha)}$. In contrast, the linear risk satisfies
$R_{\textup{lin}}^{*}(X)\gtrsim\inf_{n}\left\{n^{\beta}w_{\textup{nl}}^{2}(n)+n\sigma^{2}\right\}\asymp\sigma^{\frac{2(\alpha-\beta)}{1+\alpha-\beta}}$
as $\sigma\downarrow 0$. Then observe that $\frac{\alpha}{1+\alpha}-\frac{\delta}{1+\delta}=\frac{\alpha-\delta}{1+\alpha+\delta+\alpha\delta}$, and set $\delta=\alpha-\beta$. ∎
The role of quadratic convexity of $X$ differs between Gaussian sequence models and stochastic optimization problems, however, and the remainder of this section explores two extended examples highlighting this. We focus on sets $X\subset\mathbf{R}^{\mathbf{N}}$ in sequence space.
5.1 A constraint set requiring nonlinearity only in Gaussian sequence models
We show that for a large family of $\ell_{1}$-bodies $X$, minimax (rate) optimal estimation requires nonlinearity for Gaussian sequence models but not in stochastic optimization. Take
$X:=\bigg{\{}x\in\mathbf{R}^{\mathbf{N}}\mid\sum_{j=1}^{\infty}a_{j}|x_{j}|\leq 1\bigg{\}}~{}~{}\mbox{so}~{}~{}\mathsf{QHull}(X)=\bigg{\{}x\in\mathbf{R}^{\mathbf{N}}\mid\sum_{j=1}^{\infty}a_{j}^{2}x_{j}^{2}\leq 1\bigg{\}},$
where $(a_{j})$ is a nondecreasing positive sequence where (w.l.o.g.) we take $a_{1}=1$. Computing the $\ell_{1}$-diameters of $X$ and its quadratic hull, we then have
$\sup_{x\in X}\left\|{x}\right\|_{1}=1~{}~{}\mbox{and}~{}~{}\sup_{x\in\mathsf{QHull}(X)}\left\|{x}\right\|_{1}=\bigg{(}\sum_{j=1}^{\infty}a_{j}^{-2}\bigg{)}^{1/2}$ (17)
by the Cauchy-Schwarz inequality. On the other hand, because the coefficients $a_{j}$ are increasing, we can compute both the linear and nonlinear widths via
$0pt^{2}(n)=\sup\bigg{\{}\sum_{j=n+1}^{\infty}x_{j}^{2}\mid\sum_{j=1}^{\infty}a_{j}^{2}x_{j}^{2}\leq 1,~{}x_{1}\geq x_{2}\geq\cdots\geq 0\bigg{\}}$
and
$w_{\textup{nl}}^{2}(n)=\sup\bigg{\{}\sum_{j=n+1}^{\infty}x_{j}^{2}\mid\sum_{j=1}^{\infty}a_{j}x_{j}\leq 1,~{}x_{1}\geq x_{2}\geq\cdots\geq 0\bigg{\}}.$
By convexity (we maximize a convex function over a convex set in each case), for each width the maximizing point takes the form $(t,t,\ldots,t,0,\ldots)$, with $m$ repeated values $t$. Then we obtain
$0pt^{2}(n)$ $=\sup_{m\geq n,t\geq 0}\bigg{\{}(m-n)t^{2}\mid t^{2}\sum_{j=1}^{m}a_{j}^{2}\leq 1\bigg{\}}=\sup_{m\geq n}\frac{m-n}{\sum_{j=1}^{m}a_{j}^{2}}$ (18)
$w_{\textup{nl}}^{2}(n)$ $=\sup_{m\geq n,t\geq 0}\bigg{\{}(m-n)t^{2}\mid t\sum_{j=1}^{m}a_{j}\leq 1\bigg{\}}=\sup_{m\geq n}\frac{m-n}{(\sum_{j=1}^{m}a_{j})^{2}}.$
Comparing the $\ell_{1}$-diameters (17) and the widths (18), we see that whenever $\sum_{j}a_{j}^{-2}$ is summable Euclidean gradient methods are minimax (rate) optimal for stochastic optimization. If $a_{j}=j^{\alpha/2}$ for some $\alpha>1$, however, we have
$0pt^{2}(n)$ $=\sup_{m\geq n}\frac{m-n}{\sum_{j=1}^{m}j^{\alpha}}\asymp\sup_{m\geq n}\frac{m-n}{m^{1+\alpha}}\asymp\frac{1}{n^{\alpha}}~{}~{}\mbox{while}$
$w_{\textup{nl}}^{2}(n)$ $=\sup_{m\geq n}\frac{m-n}{(\sum_{j=1}^{m}j^{\alpha/2})^{2}}\asymp\sup_{m\geq n}\frac{m-n}{m^{2+\alpha}}\asymp\frac{1}{n^{1+\alpha}}$
[…]
Corollary 11.
Let $X=\{x\mid\sum_{j=1}^{\infty}a_{j}|x_{j}|\leq 1\}$, where $a_{j}=j^{\alpha/2}$ and $\alpha>1$. Then minimax rate optimal estimation for Gaussian sequence models requires that the estimator $\widehat{x}$ be nonlinear, while Euclidean gradient methods are minimax rate optimal for stochastic optimization.
5.2 A constraint set requiring nonlinearity only in stochastic optimization
To contrast with Corollary 11, we can also give families of underlying constraint sets $X$ for which only nonlinear methods can be rate-optimal for stochastic optimization problems, while linear estimators $\widehat{x}=Ay$ can be rate-optimal in Gaussian sequence models. At the grossest level, we construct sets $X$ for which $\sup_{x\in X}\left\|{x}\right\|_{1}\lesssim 1$ while $\sup_{x\in\mathsf{QHull}(X)}\left\|{x}\right\|_{1}=+\infty$, while $w_{\textup{nl}}^{2}(n)$ and $0pt^{2}(n)$ are comparable. To give a slightly more nuanced picture, we consider the rates at which the two $\ell_{1}$-diameters approach +$\infty$+ ∞, comparing
$\sup_{x\in X}\langle\mathbf{1}_{n},x\rangle~{}~{}\mbox{and}~{}~{}\sup_{x\in\mathsf{QHull}(X)}\langle\mathbf{1}_{n},x\rangle,$
where $\mathbf{1}_{n}\in\mathbf{R}^{\mathbf{N}}$ denotes the sequence with $1$ in its first $n$ positions and 00 elsewhere.
Here, we elaborate on the $\ell_{1}$ bodies yielding Corollary 11 to consider smaller axis-aligned polyhedra. Letting $e_{j}$ be the basis vectors (i.e. sequences with a $1$ in position $j$ and 0 elsewhere), let $a_{j}$ be a nondecreasing sequence with $a_{1}=1$ and $b_{j}$ be arbitrary (for now), define the two sets
$C_{0}:=\left\{\sigma_{j}a_{j}e_{j}\right\}_{j\in\mathbf{N}}~{}~{}~{}\mbox{and}~{}~{}~{}C_{1}:=\bigg{\{}\frac{1}{Z(n)}\sum_{j=1}^{n}\sigma_{j}b_{j}e_{j}\bigg{\}}_{n\in\mathbf{N}},~{}~{}~{}\mbox{where}~{}\sigma_{j}\in\{\pm 1\}.$
We choose the normalizing constants $Z(n)$ so that the points in $C_{1}$ all lie in
$\mathsf{QHull}(C_{0})=\mathsf{QHull}\bigg{\{}x\mid\sum_{j=1}^{\infty}a_{j}|x_{j}|\leq 1\bigg{\}}=\bigg{\{}x\mid\sum_{j=1}^{\infty}a_{j}^{2}x_{j}^{2}\leq 1\bigg{\}},$
i.e. $Z(n)=(\sum_{j=1}^{n}a_{j}^{2}b_{j}^{2})^{1/2}$ so that $Z(n)^{-2}\sum_{j=1}^{n}a_{j}^{2}b_{j}^{2}=1$. Then the set
$X:=\textup{Conv}\left(C_{0}\cup C_{1}\right)~{}~{}\mbox{satisfies}~{}~{}\mathsf{QHull}(X)=\mathsf{QHull}(C_{0}).$ (19)
We obtain the following corollary of the our convergence guarantees in Section 4 and the technical lemmas we provide in Appendix E.
Corollary 12.
Let $a_{j}=j^{\alpha/2}$ for some $10<\alpha<$ 1 and $b_{j}^{2}=2^{j}$ in the construction of $X$ above. Then
$n^{-\alpha}\leq w_{\textup{nl}}^{2}(n)\leq 0pt^{2}(n)\lesssim n^{-\alpha},$
while
$\frac{\sup_{x\in\mathsf{QHull}(X)}\langle\mathbf{1}_{n},x\rangle}{\sup_{x\in X}\langle\mathbf{1}_{n},x\rangle}\gtrsim n^{\frac{1-\alpha}{2}}.$
In particular, linear methods are rate optimal for estimation in Gaussian sequence models, while stochastic optimization over $X$ requires nonlinear methods.
Summarizing the conclusions of the corollary, as $\alpha\downarrow 0$, the ratio of the $\ell_{1}$-diameters of $\mathsf{QHull}(X)$ and $X$ grows as $\sqrt{n}$, which by Cauchy-Schwarz is as large as possible: there is a large gap between the achievable optimization performance of nonlinear methods, as Proposition 10 demonstrates, and linear methods, whose regret necessarily scales as the $\ell_{1}$-diameter of $\mathsf{QHull}(X)$ (Theorem 5). Yet the nonlinear widths are comparable for all $n$, so linear methods are minimax rate optimal as $\sigma^{2}\downarrow 0$.
6 The need for adaptive methods
We have so far demonstrated that diagonal re-scaling is sufficient to achieve minimax optimal rates for problems over quadratically convex constraint sets. In practice, however, it is often the case that we do not know the gradient geometry in advance, precluding selection of the optimal linear pre-conditioner. To address this problem [19, 26], adaptive gradient methods choose, at each step, a (usually diagonal) matrix $\Lambda_{i}$ conditional on the subgradients observed thus far, $\{g_{l}\}_{l\leq i}$. The algorithm then updates the iterate based on the distance generating function $h_{i}(x):=\frac{1}{2}x^{\top}\Lambda_{i}x$. In this section, we present a problem instance showing that when the “scale” of the subgradients varies across dimensions, adaptive gradient methods are crucial to achieve low regret. While there exists an optimal pre-conditioner, if we do not assume knowledge of the geometry in advance, AdaGrad [19] achieves the minimax optimal regret while standard (non-adaptive) subgradient methods can be $\sqrt{n}$ suboptimal on the same problem.
We consider the following setting: $X=\mathbf{B}_{\infty}(0,1)$ and $\gamma_{\beta}(g)=\left\|{\beta\odot g}\right\|_{1}$, for an arbitrary $\beta\in\mathbf{R}^{n},\beta\succ 0$. Intuitively, $\beta_{j}$ corresponds to the “scale” of the $j$-th dimension. On this problem, a straightforward optimization of the regret bound (12) guarantees that stochastic gradient methods achieve regret $\sqrt{nk}/\min_{j}\beta_{j}$. We exhibit a problem instance in the following proof of Theorem 6 such that, for any stepsize $\alpha$, online gradient descent attains this worst-case regret.
Theorem 6.
Let $\mathsf{Regret}_{k,\alpha}(x)=\sum_{i\leq k}g_{i}^{\top}(x_{i}-x)$ denote the regret of the online gradient descent method with stepsize $\alpha\geq 0$ for linear functions $F_{i}(x)=g_{i}^{\top}x$. For any choice of $\alpha\geq 0$ and $\beta\succ 0$, there exists a sequence of vectors $\{g_{i}\}_{i\leq k}\subset\mathbf{R}^{n}$, $\gamma_{\beta}(g_{i})\leq 1$ and point $x\in X$ such that
$\mathsf{Regret}_{k,\alpha}(x)\geq\frac{1}{2}\min\left\{\frac{nk}{2\left\|{\beta}\right\|_{1}},\frac{\sqrt{2nk}}{\min_{j\leq n}\beta_{j}}\right\}.$
Proof.
The proof follows similar lines as those we develop in the proof of Theorem 4, but choosing different $u,v\in\mathbf{R}^{n}$. Let $\alpha\geq 0$ be a stepsize. We consider linear functions $F_{i}(x):=g_{i}^{\top}x$ with $\left\|{\beta\odot g_{i}}\right\|_{1}\leq 1$. Let $\{x_{i}\}_{i\leq k}$ be the iterates of online gradient descent. The regret with respect to $x\in\mathbf{R}^{n}$ is
$\mathsf{Regret}_{k,\alpha}(x)=\sum_{i\leq k}g_{i}^{\top}(x_{i}-x).$
This yields
$\mathsf{Regret}_{k,\alpha}^{\ast}=\sup_{\left\|{x}\right\|_{\infty}\leq 1}\mathsf{Regret}_{k,\alpha}(x)=\bigg{\|}{\sum_{i\leq k}g_{i}}\bigg{\|}_{1}+\frac{\alpha}{2}\sum_{i\leq k}\left\|{g_{i}}\right\|_{2}^{2}-\frac{\alpha}{2}\bigg{\|}{\sum_{i\leq k}g_{i}}\bigg{\|}_{2}^{2}.$
Let $d=\arg\min_{j\leq n}\beta_{j}$, we choose
$u=e_{d}/\beta_{d}~{}~{}\mbox{and}~{}~{}v=\frac{\mathbf{1}}{\left\|{\beta}\right\|_{1}}.$
For $\delta\in[0,1]$, we now choose the vectors $g_{i}\in\mathbf{R}^{n}$ as follows:
  1. (a)
    $g_{i}=u$ for $k/4$ of the indices $i\in[k]$.
  2. (b)
    $g_{i}=-u$ for $k/4$ of the indices $i\in[k]$.
  3. (c)
    $g_{i}=v$ for $\frac{k}{4}(1+\delta)$ of the indices $i\in[k]$.
  4. (d)
    $g_{i}=-v$ for $\frac{k}{4}(1-\delta)$ of the indices $i\in[k]$.
For this construction, we lower bound the regret
$\mathsf{Regret}_{k,\alpha}^{\ast}$ $\geq\sup_{0\leq\delta\leq 1}\left\{\frac{k\delta}{2}\left\|{v}\right\|_{1}+\frac{k\alpha}{4}\left\|{u}\right\|^{2}_{2}-\frac{\alpha\delta^{2}k^{2}}{8}\left\|{v}\right\|_{2}^{2}\right\}$
$\geq\frac{k\alpha}{4}\left\|{u}\right\|_{2}^{2}+\frac{k\left\|{v}\right\|_{1}}{4}\min\left\{1,\frac{2\left\|{v}\right\|_{1}}{k\alpha\left\|{v}\right\|_{2}^{2}}\right\}.$ (20)
If the stepsize is too small (i.e. $\alpha\leq\frac{2}{k}\frac{\|{v}\|_{1}}{\|{v}\|^{2}_{2}}$) then (6) becomes
$\mathsf{Regret}_{k,\alpha}^{\ast}\geq\frac{kn}{4\left\|{\beta}\right\|_{1}}.$
In the other case that $\alpha>\frac{2}{k}\frac{\|{v}\|_{1}}{\|{v}\|^{2}_{2}}$, (6) yields
$\mathsf{Regret}_{k,\alpha}^{\ast}\geq\frac{k}{4\alpha}\left\|{u}\right\|_{2}^{2}+\frac{\left\|{v}\right\|_{1}^{2}}{\left\|{v}\right\|_{2}^{2}}\frac{\alpha}{2}\geq\frac{\sqrt{2}}{2}\frac{\sqrt{kn}}{\min_{j\leq n}\beta_{j}},$
which is the desired result. ∎
In contrast, AdaGrad [19] achieves regret $\sqrt{k}\left\|{1/\beta}\right\|_{2}$, demonstrating suboptimality gap as large as $\sqrt{n}$ for some choices of $\beta$. Indeed, let $\mathsf{Regret}_{k,\mathsf{AdaGrad}}(x)$ be the regret of AdaGrad. Then
$\mathsf{Regret}_{k,\mathsf{AdaGrad}}(x)\leq 2\sqrt{2}\sum_{j\leq n}\sqrt{\sum_{i\leq k}g_{i,j}^{2}}.$
(see Duchi et al. [19, Corollary 6]), and by Cauchy-Schwarz,
$\sum_{j\leq n}\sqrt{\sum_{i\leq k}g_{i,j}^{2}}=\sum_{j\leq n}\frac{1}{\beta_{j}}\sqrt{\sum_{i\leq k}\beta_{j}^{2}g_{i,j}^{2}}\leq\left\|{1/\beta}\right\|_{2}\sqrt{\sum_{i\leq k}\left\|{\beta\odot g_{i}}\right\|^{2}_{2}}\leq\sqrt{k}\left\|{1/\beta}\right\|_{2}.$
To concretely consider different scales across dimensions, take $\beta_{j}=j$. Theorem 6 guarantees that there exists a collection of linear functions such that stochastic gradient methods suffer regret $\Omega(1)\sqrt{nk}$. Given that $\|{1/\beta}\|_{2}\leq\sqrt{\zeta(2)}\leq\pi/\sqrt{6}$, AdaGrad achieves regret $O(1)\sqrt{k}$—amounting to a suboptimality gap of order $\sqrt{n}$—exhibiting the need for adaptivity. This $\sqrt{n}$ gap is also the largest possible over subgradient methods, which achieve regret $\sqrt{n\sum_{i\leq k}\left\|{g_{i}}\right\|_{2}^{2}}\leq\sqrt{n}\sum_{j\leq n}\sqrt{\sum_{i\leq k}g_{i,j}^{2}}$ for $X=\mathbf{B}_{\infty}(0,1)$. Finally, we note in passing that AdaGrad is minimax optimal on this class of problems via a straightforward application of Theorem 1.
7 Discussion
We provide concrete foundations to compare adaptive, mirror, or standard gradient methods, showing how problem geometry necessarily impacts convergence. This paper puts a particular computational spin on optimization by connecting to Gaussian sequence models and linear versus nonlinear updates, which we advocate for its ability to paint a different picture than pure polynomial versus non-polynomial computational complexity. This perspective draws from information-based models in optimization [27, 1] and models in scientific computing where one uses certain families of operations—e.g., matrix vector multiplies—to build up optimal algorithms under these constraints [39, 3]. We hope to see more exploration in these directions.
There remain a few additional open questions. The lower bound in Corollary 8 exhibiting a scaling of $\sup_{x\in X}\sup_{\gamma(g)\leq 1}x^{\top}g$, is sharp in some cases, but not in others. A natural conjecture that we leave open for now is that this bound is tight whenever the set $G=\{g\mid\gamma(g)\leq 1\}$ is quadratically convex. We believe this to be true because of the connections between martingale type and the regret guarantees of mirror descent [38]: in brief, whenever the norm $\gamma$ satisfies a so-called type-2 inequality [32], meaning roughly that if $S_{n}=\sum_{i=1}^{n}g_{i}$ for a martingale sequence of $g_{i}\in G$, then $\mathbf{E}[\gamma(S_{n})^{2}]\leq C\sum_{i=1}^{n}\mathbf{E}[\gamma(g_{i})^{2}]$ for a reasonable constant $C$. All the “usual” quadratically convex norms—e.g., $\ell_{q}$ norms for $q\geq 2$—satisfy such inequalities [21], and so it is natural to imagine that quadratic convexity coincides with such inequalities. We leave this open.
While Section 6 emphasizes the importance of adaptivity, the picture is not fully complete: for example, in the case of quadratically convex constraint sets, while the best diagonal pre-conditioner achieves optimal rates, the extent to which adaptive gradient algorithms find this optimal pre-conditioner remains an open question. Another avenue to explore involves the many flavors of adaptivity—while the minimax framework assumes knowledge of the problem setting (e.g. a bound on the domain or the gradient norms), it is often the case that such parameters are unknown to the practitioner. To what extent can adaptivity mitigate this and achieve optimal rates, and is minimax (i.e. worst-case) optimality truly the right measure of performance?
Acknowledgments.
We thank Aditya Grover, Annie Marsden, and Hongseok Namkoong for valuable comments, and Quentin Guignard for pointing us to the Banach-Mazur distance for Theorem 4.
Appendix A Duality results and minimax linear estimators
In this section, we collect the deferred proofs of various optimality results for the Gaussian sequence model in Section 2.1 that we use.
A.1 Proof of Proposition 1
Recalling the form (6) of the risk $R(A,x)=x^{T}(A-I)^{T}(A-I)x+\sigma^{2}\left\|{A}\right\|_{\rm Fr}^{2}$, it is immediate that $A$ is unique, as $\left\|{A}\right\|_{\rm Fr}^{2}$ is strongly convex in $A$ and $\sup_{x\in X}\left\|{(A-I)x}\right\|_{2}^{2}$ is convex in $A$. As $X$ is compact, the risk $R^{*}(A,X)$ is finite for all $A$, and hence its minimum is attained.
We now consider the saddle point problem
$\inf_{A}\sup_{x\in X}R(A,x)=\inf_{A}\sup_{x\in X}\left\{x^{T}A^{T}Ax-2x^{T}Ax+\sigma^{2}\left\|{A}\right\|_{\rm Fr}^{2}\right\}.$
By lifting to the space $\mathcal{P}(X)$ of probability measures defined on $X$, we have for each $A$ that
[…]
We shall apply Fan’s minimax theorem [22], a specialization of Sion’s minimax theorem [37] to construct the $A$ minimizing this worst-case risk.
Lemma 5.
Let $X$ and $Y$ be compact convex subsets of (possibly distinct) topological vector spaces and $L:X\times Y\to\mathbf{R}$ be convex in its first argument, concave in its second, and continuous. Then
$\inf_{x\in X}\sup_{y\in Y}L(x,y)=\sup_{y\in Y}\inf_{x\in X}L(x,y)$
and the infimum and supremum are attained.
[…]
$W(\mu,\nu)=\sup_{\left\|{f}\right\|_{\textup{Lip}}\leq 1}\int f(d\mu-d\nu).$
The Wasserstein distance metrizes convergence in distribution $\mu_{n}\stackrel{{\scriptstyle d}}{{\to}}\mu$ if the measures $\mu_{n},\mu$ have bounded first moment and turns the space of measures on $X$ into a (topological) vector space.
In our case, we define
[…]
which is clearly convex/concave and continuous in $A$. To see the continuity of $L$ in $\nu$, we note that for any matrix $B$ and any $x,y\in X$ we have
$\langle xx^{T}-yy^{T},B\rangle=(x-y)^{T}(B+B^{T})(x+y)/2\leq 2\left|\!\left|\!\left|{B}\right|\!\right|\!\right|_{\rm op}\left\|{x-y}\right\|_{2}\textup{diam}(X),$
so that $x\mapsto\langle xx^{T},B\rangle$ is Lipschitz and
$|L(A,\nu)-L(A,\mu)|\leq 2\left|\!\left|\!\left|{A-I}\right|\!\right|\!\right|_{\rm op}^{2}\textup{diam}(X)\cdot W(\nu,\mu).$
[…]
$\inf_{B}L(B,\nu)=L(A,\nu)={\rm tr}((A-I)^{T}(A-I)X_{\nu})+\sigma^{2}\left\|{A}\right\|_{\rm Fr}^{2}=\sup_{\mu\in\mathcal{P}(X)}L(A,\mu).$ (21)
[…]
$AX_{\nu}+X_{\nu}A^{T}-2X_{\nu}+\sigma^{2}(A+A^{T})=0.$
If $X_{\nu}$ has spectral decomposition $X_{\nu}=U\Lambda U^{T}$ we let $A=UDU^{T}$ for a diagonal matrix $D$ to be determined, and it is evidently enough to solve
$2D\Lambda-2\Lambda+2\sigma^{2}D=0,~{}~{}~{}\mbox{or}~{}~{}~{}D=(\Lambda+\sigma^{2}I)^{-1}\Lambda.$
In particular, the choice $A=(X_{\nu}+\sigma^{2}I)^{-1/2}X_{\nu}(X_{\nu}+\sigma^{2}I)^{-1/2}$ is optimal; it is also unique for the given $\nu$ as $A\mapsto L(A,\nu)$ is strongly convex in $A$.
Finally, we show that without loss of generality, we may take $A$ to be diagonal. If $S$ is a diagonal matrix of independent random signs, then $\mathbf{E}_{\nu}[(Sx)(Sx)^{T}]=\textup{diag}(X_{\nu})=\mathbf{E}_{\nu}[\textup{diag}(x)^{2}]$. Let $\overline{\nu}$ be the measure on $X$ induced by drawing $x\sim\nu$ and then multiplying $x$ by the random signs $Sx$. Notably, we have ${\rm tr}(X_{\nu})={\rm tr}(X_{\overline{\nu}})$ and ${\rm tr}(DX_{\nu})={\rm tr}(DX_{\overline{\nu}})$ for any diagonal matrix $D$, as $\textup{diag}(X_{\nu})=\textup{diag}(X_{\overline{\nu}})$. Suppose for the sake of contradiction that $L(\textup{diag}(A),\nu)>L(A,\nu)$. In this case, $A$ must be non-diagonal, and so we have
$L(\textup{diag}(A),\nu)$ $={\rm tr}(\textup{diag}(A)^{2}X_{\nu})-2{\rm tr}(\textup{diag}(A)X_{\nu})+{\rm tr}(X_{\nu})+\sigma^{2}\left\|{\textup{diag}(A)}\right\|_{\rm Fr}^{2}$
$={\rm tr}(\textup{diag}(A)^{2}X_{\overline{\nu}})-2{\rm tr}(\textup{diag}(A)X_{\overline{\nu}})+{\rm tr}(X_{\overline{\nu}})+\sigma^{2}\left\|{\textup{diag}(A)}\right\|_{\rm Fr}^{2}$
<$\stackrel{{\scriptstyle(\star)}}{{<}}{\rm tr}(\textup{diag}(A^{2})X_{\overline{\nu}})-2{\rm tr}(\textup{diag}(A)X_{\overline{\nu}})+{\rm tr}(X_{\overline{\nu}})+\sigma^{2}\left\|{A}\right\|_{\rm Fr}^{2}$
$={\rm tr}(A^{2}X_{\overline{\nu}})-2{\rm tr}(\textup{diag}(A)X_{\overline{\nu}})+{\rm tr}(X_{\overline{\nu}})+\sigma^{2}\left\|{A}\right\|_{\rm Fr}^{2},$
where inequality ()($\star$)( ⋆ ) follows because $\left\|{A}\right\|_{\rm Fr}>\left\|{\textup{diag}(A)}\right\|_{\rm Fr}$ while $\textup{diag}(A)^{2}\preceq\textup{diag}(A^{2})$. The final line follows because $X_{\overline{\nu}}$ is diagonal, so $\textup{diag}(A^{2}X_{\overline{\nu}})=\textup{diag}(A^{2})X_{\overline{\nu}}$. Finally, noting that ${\rm tr}(\textup{diag}(A)X_{\overline{\nu}})={\rm tr}(AX_{\overline{\nu}})$, we see that $L(\textup{diag}(A),\nu)<L(A,\overline{\nu})$, and so we have demonstrated that $L(A,\nu)<L(A,\overline{\nu})$. But this contradicts the assumed maximality of $\nu$, and so it must be the case that $A$ is diagonal.
Now that we have $A$ diagonal, the claimed equality is immediate, and we also notice that $L(D,\nu)=L(D,\overline{\nu})$ for any diagonal $D$.
A.2 Proof of Corollary 1
Proposition 1 implies that
$\inf_{A}\sup_{x\in X}\mathbf{E}_{x}\left[\|{Ay-x}\|_{2}^{2}\right]$ $=\inf_{d\in\mathbf{R}^{n}}\sup_{x\in X}\sum_{j=1}^{n}\left((d_{j}-1)^{2}x_{j}^{2}+\sigma^{2}d_{j}^{2}\right)$
$=\inf_{d}\sup_{v\in\mathsf{SqHull}(X)}\sum_{j=1}^{n}\left((d_{j}-1)^{2}v_{j}+\sigma^{2}d_{j}^{2}\right)$
$=\sup_{v\in\mathsf{SqHull}(X)}\inf_{d}\left\{\sum_{j=1}^{n}\left((d_{j}-1)^{2}v_{j}+\sigma^{2}d_{j}^{2}\right)\right\},$ (22)
where equality (22) is a standard convex/concave saddle-point result (we may without loss of generality restrict $d$ to the set $[0,1]^{n}$). Continuing the equalities, we have
$\inf_{d}\left\{(d_{j}-1)^{2}v_{j}+\sigma^{2}d_{j}^{2}\right\}=\frac{\sigma^{2}v_{j}}{v_{j}+\sigma^{2}},$
which implies the corollary.
A.3 Proof of Corollary 2
For any linear operator $A:\mathbf{R}^{\mathbf{N}}\to\mathbf{R}^{\mathbf{N}}$, we may write
$(Ay)_{j}=a_{j}(y)=a_{j}(x+\xi)=a_{j}(x)+a_{j}(\xi)$
for each $j$. Let $\Pi_{n}:\mathbf{R}^{\mathbf{N}}\to\mathbf{R}^{n}$ be the projection onto the first $n$ coordinates of a vector and $Z_{n}:\mathbf{R}^{\mathbf{N}}\to\mathbf{R}^{\mathbf{N}}$ be the projection zeroing the first $n$ elements. Then
$\inf_{A}\sup_{x\in X}\mathbf{E}_{x}\left[\|{Ay-x}\|_{2}^{2}\right]\geq\inf_{A}\sup_{x\in\Pi_{n}X}\mathbf{E}_{x}[\left\|{Ay-x}\right\|_{2}^{2}]=\inf_{A}\sup_{x\in\Pi_{n}X}\sum_{j=1}^{n}\mathbf{E}_{x}[(a_{j}((x,\mathbf{0})+\xi)-x_{j})^{2}],$
where $(x,\mathbf{0})\in\Pi_{n}X\times\mathbf{R}^{\mathbf{N}}$. Then by linearity, for $x\in\mathbf{R}^{n}$ we write
$a_{j}((x,\mathbf{0})+\xi)=\varphi_{j}(x)+a_{j}(Z_{n}\xi)+a_{j}((I-Z_{n})\xi)$
where $\varphi_{j}:\mathbf{R}^{n}\to\mathbf{R}$ is the linear function $\varphi_{j}(x)=a_{j}((x,\mathbf{0}))$. But $Z_{n}\xi$ and $(I-Z_{n})\xi=(\xi_{1},\ldots,\xi_{n},\mathbf{0})$ are independent, and because $\mathbf{E}[a_{j}(Z_{n}\xi)]=0$ we obtain
$\mathbf{E}_{x}[(a_{j}((x,\mathbf{0})+\xi)-x_{j})^{2}]$ $=\mathbf{E}[(\varphi_{j}(x+[\xi_{j}]_{j\leq n})-x_{j}+a_{j}(Z_{n}\xi))^{2}]$
$=\mathbf{E}[(\varphi_{j}(x+[\xi_{j}]_{j\leq n})-x_{j})^{2}]+\mathop{\rm Var}(a_{j}(Z_{n}\xi)).$
The optimal choice of $a_{j}$ then necessarily satisfies $a_{j}(Z_{n}u)=0$ for all $u$. Thus, by restricting to finite dimensions, we use Proposition 1 to see that for any $n$,
$\inf_{A}\sup_{x\in X}\mathbf{E}_{x}\left[\|{Ay-x}\|_{2}^{2}\right]$ $\geq\inf_{d\in\mathbf{R}^{n}}\sup_{x\in X}\sum_{j=1}^{n}\left((d_{j}-1)^{2}x_{j}^{2}+\sigma^{2}d_{j}^{2}\right)$
$=\sup_{v\in\mathsf{SqHull}(X)}\inf_{d}\sum_{j=1}^{n}\left((d_{j}-1)^{2}v_{j}+\sigma^{2}d_{j}^{2}\right)=\sup_{v\in\mathsf{SqHull}(X)}\sum_{j=1}^{n}\frac{\sigma^{2}v_{j}}{v_{j}+\sigma^{2}}$
as in the proof of Corollary 1.
By compactness, for each $\epsilon>0$, we can choose $N<\infty$ such that $\sup_{x\in X}\sum_{j>N}x_{j}^{2}<\epsilon$. We thus have upper bound
$\inf_{A}R^{*}(A,X)\leq\inf_{A}\sup_{x\in\Pi_{N}X}\mathbf{E}_{x}[\left\|{Ay-x}\right\|_{2}^{2}]+\epsilon.$
Apply the same proof technique as that in Corollary 1 to obtain
$\inf_{A}R^{*}(A,\Pi_{N}X)=\sup_{v\in\mathsf{SqHull}(\Pi_{N}X)}\sum_{j=1}^{N}\frac{\sigma^{2}v_{j}}{v_{j}+\sigma^{2}}=\sup_{v\in\mathsf{SqHull}(X)}\sum_{j=1}^{N}\frac{\sigma^{2}v_{j}}{v_{j}+\sigma^{2}}.$
Now use that $\epsilon>0$ was arbitrary, $X$ is compact, and take $n$ and $N$ to infinity.
A.4 Proof of Proposition 2
We apply Le Cam’s two point method via the reduction from estimation to testing that Lemma 1 implies. Consider for any fixed $b>0$ the problem of estimating a single value $x\in[-b,b]$ with $y\sim\mathsf{N}(x,\sigma^{2})$. Then for $v\in\{-1,1\}$, letting $P_{v}$ be the normal distribution $\mathsf{N}(\delta v,\sigma^{2})$ for some $\delta\in[0,b]$ to be chosen, we have
$\inf_{\widehat{x}}\sup_{x\in[-b,b]}\mathbf{E}_{x}[(\widehat{x}-x)^{2}]\geq\inf_{\widehat{x}}\frac{1}{2}\left\{\mathbf{E}_{P_{1}}[(\widehat{x}-x)^{2}]+\mathbf{E}_{P_{-1}}[(\widehat{x}-x)^{2}]\right\}\geq\frac{\delta^{2}}{2}\left(1-\left\|{P_{-1}-P_{1}}\right\|_{\mathsf{tv}}\right).$
Recalling the Hellinger distance $d_{\textup{hel}}^{2}(P,Q)=1-\int\sqrt{dPdQ}$ between probabilities and the standard relationship $\left\|{P_{-1}-P_{1}}\right\|_{\mathsf{tv}}\leq d_{\textup{hel}}(P_{1},P_{-1})\sqrt{2-d_{\textup{hel}}^{2}(P_{1},P_{-1})}$, we note that $d_{\textup{hel}}^{2}(\mathsf{N}(\mu_{0},\sigma^{2}),\mathsf{N}(\mu_{1},\sigma^{2}))=1-\exp(-\frac{1}{8\sigma^{2}}(\mu_{0}-\mu_{1})^{2})$ to obtain
$\inf_{\widehat{x}}\sup_{x\in[-b,b]}\mathbf{E}_{x}\left[(\widehat{x}-x)^{2}\right]\geq\sup_{0\leq\delta\leq b}\frac{\delta^{2}}{2}\left(1-\sqrt{1-\exp(-\delta^{2}/\sigma^{2})}\right)\geq\frac{\sigma^{2}\wedge b^{2}}{10}$
via the choice $\delta=\min\{\sigma,b\}$. We thus obtain for any hypercube $H=[-x_{j},x_{j}]_{j\geq 1}$ that
$R^{*}(X)\geq R^{*}(H)\geq\frac{1}{10}\sum_{j\geq 1}\sigma^{2}\wedge x_{j}^{2}.$
A.5 Proof of Corollary 6
We adapt the arguments in [17, pg. 1430]. For $N=N(\sigma,X)$ we have
$\mathbf{E}[\|{\widehat{x}-x}\|_{2}^{2}]$ $=\sum_{j=1}^{N}\mathbf{E}[(\mathsf{S}_{\lambda}(y_{j})-x_{j})^{2}]+\sum_{j>N}x_{j}^{2}$
$\leq\sigma^{2}+(1+2\log N)\sum_{j=1}^{N}x_{j}^{2}\wedge\sigma^{2}+\sum_{j>N}x_{j}^{2}\wedge\sigma^{2},$
as $x_{j}^{2}\leq\sigma^{2}$ for $j>N(\sigma,X)$, which implies the first result. Proposition 2 implies the second.
Appendix B Proofs for Section 3.2
B.1 Proof of Theorem 2
The upper bound is simply Corollary 7. For the lower bound, similar to our warm-up in Section 3.1, we consider “sparse” gradients, though instead of using Fano’s method we use Assouad’s method to more carefully relate the geometry of the norm $\gamma$ and constraint set $X$.
Let $a\in\mathbf{R}^{n}_{+}$ be such that $\mathsf{Rec}(a)\subset X$. We consider the sample space $\mathcal{S}:=\{\pm e_{j}\}_{j\leq n}$ and functions
$F(x,s):=\sum_{j\leq n}\frac{1}{\gamma(e_{j})}|s_{j}||x_{j}-a_{j}s_{j}|.$
For any $s\in\mathcal{S}$, the subdifferential $\partial_{x}F(x,s)$ has at most one non-zero coordinate; the orthosymmetry of $\gamma$ implies $F\in\mathcal{F}^{\gamma,1}$. Let $p\in\mathbf{R}^{n}_{+}$ (to be specified presently) be such that $\mathbf{1}^{\top}p=1$ and for $1\leq j\leq n$, let $\delta_{j}\in[0,1/2]$. We define the distributions $P_{v}$ on $\mathcal{S}$ by
[…]
With this choice, we evidently have
$f_{v}(x)=\mathbf{E}_{S\sim P_{v}}F(x,S)=\sum_{j\leq n}\frac{p_{j}}{\gamma(e_{j})}\left[\frac{1+\delta_{j}}{2}|x_{j}-a_{j}v_{j}|+\frac{1-\delta_{j}}{2}|x_{j}+a_{j}v_{j}|\right]$
and immediately that $\inf_{X}f_{v}=\sum_{j\leq n}\frac{p_{j}a_{j}}{\gamma(e_{j})}(1-\delta_{j})$. As a consequence, we have the Hamming separation (recall Eq. (10))
$f_{v}(x)-\inf_{X}f_{v}=\sum_{j\leq n}\frac{p_{j}a_{j}\delta_{j}}{\gamma(e_{j})}\mathbf{1}_{\mathop{\rm sign}(x_{j})\neq v_{j}},$
which allows us to apply Assouad’s method via Lemma 4.
Using the same notation as Lemma 4, we have
+$\left\|{\mathbb{P}^{k}_{+j}-\mathbb{P}^{k}_{-j}}\right\|_{\mathsf{tv}}^{2}\leq\frac{1}{2}D_{\textup{kl}}\left({\mathbb{P}^{k}_{+j}}|\!|{\mathbb{P}^{k}_{-j}}\right)\leq\log 3\cdot kp_{j}\delta_{j}^{2}.$
Choosing $\delta_{j}=\min\{\frac{1}{2},\frac{1}{2\sqrt{kp_{j}\log(3)}}\}$ yields the lower bound
$\mathfrak{M}_{k}^{\mathsf{S}}(X,\gamma)\geq\frac{1}{8}\sum_{j\leq n}\frac{a_{j}}{\gamma(e_{j})}\min\left\{p_{j},\frac{\sqrt{p_{j}}}{\sqrt{k\log 3}}\right\},$
and by taking $p_{j}=(\frac{a_{j}}{\gamma(e_{j})})^{2}/\left\|{\textup{res}(a,\gamma)}\right\|_{2}^{2}$, we obtain for any $a\in X$ that
$\mathfrak{M}_{k}^{\mathsf{S}}(X,\gamma)$ $\geq\mathfrak{M}_{k}^{\mathsf{S}}(\mathsf{Rec}(a),\gamma)$
$\geq\frac{1}{8}\sum_{j\leq n}\frac{a_{j}}{\gamma(e_{j})}\min\left\{\frac{a_{j}^{2}}{\gamma(e_{j})^{2}\left\|{\textup{res}(a,\gamma)}\right\|_{2}^{2}},\frac{1}{\sqrt{k\log 3}}\frac{a_{j}}{\gamma(e_{j})\left\|{\textup{res}(a,\gamma)}\right\|_{2}}\right\}$
$=\frac{1}{8\left\|{\textup{res}(a,\gamma)}\right\|_{2}^{2}}\sum_{j=1}^{n}\frac{a_{j}^{2}}{\gamma(e_{j})^{2}}\min\left\{\frac{a_{j}}{\gamma(e_{j})},\frac{\left\|{\textup{res}(a,\gamma)}\right\|_{2}}{\sqrt{k\log 3}}\right\}$
For notational simplicity, define the set $T:=\{\textup{res}(x,\gamma))\mid x\in X\}$, which is evidently orthosymmetric and convex (it is a diagonal scaling of $X$). Then
$\mathfrak{M}_{k}^{\mathsf{S}}(X,\gamma)\geq\sup_{u\in T}\frac{1}{8\left\|{u}\right\|_{2}^{2}}\sum_{j=1}^{n}u_{j}^{2}\min\left\{u_{j},\frac{\left\|{u}\right\|_{2}}{\sqrt{k\log 3}}\right\}.$ (23)
For any vector $u\in\mathbf{R}_{+}^{n}$ and $c<1$, if we define $J=\{j\in[n]\mid u_{j}\geq\frac{c}{\sqrt{n}}\left\|{u}\right\|_{2}\}$, then
$\left\|{u}\right\|_{2}^{2}=\left\|{u_{J}}\right\|_{2}^{2}+\left\|{u_{J^{c}}}\right\|_{2}^{2}\leq\left\|{u_{J}}\right\|_{2}^{2}+\left\|{u}\right\|_{2}^{2}\sum_{j\in J^{c}}\frac{c^{2}}{n}\leq\left\|{u_{J}}\right\|_{2}^{2}+c^{2}\left\|{u}\right\|_{2}^{2},$
i.e. $\left\|{u_{J}}\right\|_{2}\geq\sqrt{1-c^{2}}\left\|{u}\right\|_{2}$. Now, fix $d\in\mathbf{N}$. If in the supremum (23) we consider any vector $u\in T,u\geq 0$ satisfying $\left\|{u}\right\|_{0}\leq d$, then setting the index set $J=\{j:u_{j}\geq\left\|{u}\right\|_{2}/\sqrt{k\log 3}\}=\{j:u_{j}\geq\left\|{u}\right\|_{2}/\sqrt{d(k/d)\log 3}\}$ we have
$\mathfrak{M}_{k}^{\mathsf{S}}(X,\gamma)$ $\geq\frac{1}{8\left\|{u}\right\|_{2}^{2}}\sum_{j=1}^{n}u_{j}^{2}\min\left\{u_{j},\frac{\left\|{u}\right\|_{2}}{\sqrt{k\log 3}}\right\}\geq\frac{1}{8\left\|{u}\right\|_{2}^{2}}\sum_{j\in J}u_{j}^{2}\frac{\left\|{u}\right\|_{2}}{\sqrt{k\log 3}}$
$\geq\frac{1}{8}\left(1-\frac{d}{k\log 3}\right)\frac{\left\|{u}\right\|_{2}}{\sqrt{k\log 3}}.$
Taking a supremum over $u$ with $\left\|{u}\right\|_{0}\leq d$ gives the theorem.
B.2 Proof of Corollary 9
Given the proof of Theorem 2, the proof is nearly immediate. Let $p\in[1,2]$, $\beta\in\mathbf{R}_{++}^{n}$, and $\gamma(v)=\left\|{\beta\odot v}\right\|_{p}$. For the lower bound, the final display of the proof of Theorem 2 above guarantees the lower bound $\mathfrak{M}_{k}^{\mathsf{S}}(X,\gamma)\geq\frac{1}{16}\left\|{u}\right\|_{2}/\sqrt{k}$ for all $u\in\{\textup{res}(x,\gamma)\mid x\in X\}$ and $k\geq 2n$. We first observe that $\mathsf{QHull}(\mathbf{B}_{\gamma}(0,1))=\{v\in\mathbf{R}^{n}\mid\left\|{\beta\odot v}\right\|_{2}\leq 1\}$. Thus, the upper bound in Theorem 2 is
$\mathfrak{M}_{k}^{\mathsf{R}}(X,\gamma)\leq\frac{1}{\sqrt{k}}\sup_{x\in X}\sup_{g:\left\|{\beta\odot g}\right\|_{2}\leq 1}x^{\top}g.$
Using
$\sup_{g:\left\|{\beta\odot g}\right\|_{2}\leq 1}u^{\top}g=\sup_{z:\left\|{z}\right\|_{2}\leq 1}u^{\top}\left(z/\beta\right)=\left\|{u/\beta}\right\|_{2},$
and recalling $\beta_{j}=\gamma(e_{j})$ concludes the proof.
Appendix C Proofs for Section 4
C.1 Proof of Theorem 3
Let us tackle the first case stated in the theorem; we reduce the second case to the first one by scaling the dimension. As in the proofs of Proposition 7 and 8, let $c_{p}=n^{-1/p}$, and let $p^{*}=\frac{p}{p-1}$ be conjugate to $p$, so that $c_{p^{*}}\left\|{\mathbf{1}}\right\|_{p^{*}}=1$.
C.1.1 Case $1\leq p\leq 1+1/\log(2n)$
Lemma 3 always implies the lower bound $1/\sqrt{k}$, which we may obtain by reducing to a lower-dimensional problem, so we assume without loss of generality that $n\geq 8$.
Separation
Take $\mathcal{V}=\{\pm e_{j}\}_{j\leq n}$, and for $v=\pm e_{j}\in\mathcal{V}$, we define $P_{v}$ on $S\in\{\pm 1\}^{n}$ by choosing coordinates of $S$ independently via
$S_{j}=\begin{cases}1&\mbox{~{}~{}with probability~{}~{}}\frac{1+\delta v_{j}}{2}\\ -1&\mbox{~{}~{}with probability~{}~{}}\frac{1-\delta v_{j}}{2},\end{cases}$
whence evidently $\mathbf{E}_{P_{v}}[S]=\delta v$. For $s\in\{\pm 1\}^{n}$, construct $F\in\mathcal{F}^{\gamma,1}$ by defining
$F(x,s):=\frac{c_{p^{*}}}{2}\left[s^{\top}x+\left({\left\|{x}\right\|_{1}-1}\right)_{+}\right].$
Then the population objective $f_{v}(x):=\mathbf{E}_{P_{v}}[F(x;S)]=\frac{c_{p^{*}}}{2}[\delta v^{\top}x+\left({\left\|{x}\right\|_{1}-1}\right)_{+}]$, which has minimizer $x_{v}=-v\in X$ and minimal value $\inf_{x}f_{v}(x)=\inf_{x\in X}f_{v}(x)=-\frac{c_{p^{*}}}{2}\delta\left\|{v}\right\|_{2}^{2}=-\frac{c_{p^{*}}}{2}\delta$. The sum $f_{v}(x)+f_{v^{\prime}}(x)=\frac{c_{p^{*}}}{2}[\delta(v+v^{\prime})^{\top}x+2\left({\left\|{x}\right\|_{1}-1}\right)_{+}]$ has minimizer $x=-\frac{1}{2}(v+v^{\prime})$. Combining these calculations gives separation
$\mathsf{d}_{\textup{opt}}(v,v^{\prime},X)=c_{p^{*}}\left[\delta-\frac{\delta}{4}\left\|{v+v^{\prime}}\right\|_{2}^{2}\right]\geq\frac{c_{p^{*}}\delta}{2}=\frac{n^{-1/p^{*}}\delta}{2},$
the inequality holding whenever $v\neq v^{\prime}\in\{\pm e_{j}\}_{j=1}^{n}$. Lemma 2 yields
$\mathfrak{M}_{k}^{\mathsf{S}}(X,\gamma)\geq\frac{\delta n^{-1/p^{\ast}}}{4}\inf_{\psi:\mathcal{S}^{k}\to\mathcal{V}}\mathbb{P}(\psi(S_{1}^{k})\neq V).$
It now remains to bound the testing error.
Bounding the testing error
Noting that $|\mathcal{V}|=\log(2n)$, we lower bound the testing error via Fano’s inequality
$\inf_{\psi:\mathcal{S}^{k}\to\mathcal{V}}\mathbb{P}(\psi(S_{1}^{k})\neq V)\geq\left(1-\frac{\mathsf{I}(S_{1}^{k};V)+\log 2}{\log(2n)}\right).$
For any $v\neq v^{\prime}\in\mathcal{V}$, we have for $\delta\in[0,\frac{1}{2}]$ that
$D_{\textup{kl}}\left({P_{v}}|\!|{P_{v^{\prime}}}\right)=\delta\log\frac{1+\delta}{1-\delta}\leq 3\delta^{2}.$
We can thus bound the mutual information between $S_{1}^{n}$ and $V$
$\mathsf{I}(S_{1}^{k};V)\leq k\max_{v\neq v^{\prime}}D_{\textup{kl}}\left({P_{v}}|\!|{P_{v^{\prime}}}\right)\leq 3k\delta^{2}.$
In the case that $n<8$, the lower bound holds trivially via Lemma 3. In the case that $n\geq 8$, choosing $\delta^{2}=\frac{\log(2n)}{6k}\wedge\frac{1}{2}$ yields
$\mathfrak{M}_{k}^{\mathsf{S}}(X,\gamma)\geq\frac{n^{-1/p^{*}}}{4}\min\left\{\sqrt{\frac{\log(2n)}{6k}},\frac{1}{2}\right\}\left(1-\frac{1}{2}-\frac{1}{4}\right),$ (24)
which is valid for all $p\in[1,2]$. In the case that $1\leq p\leq 1+1/\log(2n)$, we note that $n^{-1/p^{\ast}}=1/n^{\frac{p-1}{p}}\geq 1/e$, which yields
$\mathfrak{M}_{k}^{\mathsf{S}}(X,\gamma)\geq c\cdot\sqrt{\frac{\log(2n)}{k}}\wedge 1$
for a numerical constant $c>0$.
To conclude, we see that the upper bound follows immediately from Proposition 5. Choosing $a=1+1/\log(2n)$ and $a^{*}=\frac{a}{a-1}=1+\log(2n)$, the quantity $\frac{\sup_{x\in X}\left\|{x}\right\|_{a}\sup_{g\in\mathbf{B}_{\gamma}(0,1)}\left\|{g}\right\|_{a^{*}}}{\sqrt{a-1}\sqrt{k}}$ upper bounds the minimax regret. As $a\geq p$, $\sup_{x\in X}\|x\|_{a}=1$, and $a^{*}\leq p^{*}$ so
$\|g\|_{a^{\ast}}\leq n^{\frac{1}{a^{\ast}}-\frac{1}{p^{\ast}}}\|g\|_{p^{\ast}}\leq n^{\frac{1}{a^{\ast}}}$
as $g\in\mathbf{B}_{p^{\ast}}(0,1)$. Once we note that both $n^{1/a^{\ast}}=\exp(\frac{\log n}{\log(2n)+1})\leq e$ and $1/\sqrt{a-1}=\sqrt{\log(2n)}$, we conclude this case.
C.1.2 Case $1+1/\log(2n)<p\leq 2$
Let $n_{0}\leq n$. We can embed a function $F_{n_{0}}:\mathbf{R}^{n_{0}}\times\mathcal{S}\to\mathbf{R}$ as a function $F:\mathbf{R}^{n}\times\mathcal{S}\to\mathbf{R}$ by letting $\pi_{n_{0}}$ denote the projection onto the first $n_{0}$-components, and defining
$F(x,s)=F_{n_{0}}(\pi_{n_{0}}x,s).$
If the subgradients of $F_{n_{0}}$ lie in $\mathbf{B}_{p^{\ast}}(0,1)$, so do those of $F$. Similarly, if $x_{0}\in\{\tau\in\mathbf{R}^{n_{0}},\|\tau\|_{p}\leq 1\}$ then $x=(x_{0},\mathbf{0}_{n_{0}+1:n})\in\mathbf{B}_{p}(0,1)$. As such, any lower bound for the $n_{0}$-dimensional problem implies an identical one for all $n\geq n_{0}$-dimensional problems. For $1+1/\log(2n)<p\leq 2$, let us define $n_{0}=\lceil\exp(\frac{1}{p-1})/2\rceil$, so $n_{0}\leq n$ as desired. In the case that $p>1+1/\log 16$, Lemma 3 yields the desired lower bound. In the case that $p\leq 1+1/\log 16$, we have that $n_{0}\geq 8$, and the lower bound (24) holds so that for a numerical constant $c>0$,
$\mathfrak{M}_{k}^{\mathsf{S}}(X,\gamma)\geq cn_{0}^{-1/p^{\ast}}\cdot\sqrt{\frac{\log(2n_{0})}{n}}\wedge 1.$
Once we use that $n_{0}^{-1/p^{\ast}}\geq(1/2)^{\frac{1}{p}-1}\exp(-1/p)\geq\sqrt{2/e}$, substituting $n_{0}=\lceil\exp(\frac{1}{p-1})/2\rceil$ above gives the final lower bound
$\mathfrak{M}_{k}^{\mathsf{S}}(X,\gamma)\geq c\cdot\frac{1}{\sqrt{2(p-1)n}}\wedge 1.$
Proposition 5 yields the upper bound and concludes this proof.
C.2 Proof of Observation 4.1
The first claim of the observation is trivial. For the second and third, we use a bit of additional notation, saying that $X$ is $\beta$-self-similar if
$\sup_{x\in X}\sum_{j=1}^{\infty}j^{\beta}|x_{j}|\leq e\sup_{x\in X}\left\|{x}\right\|_{1}.$ (25)
We show that if $\lim_{\beta\downarrow 0}\sup_{x\in X}\sum_{j=1}^{\infty}j^{\beta}|x_{j}|<\infty$, then there is some $\beta>0$ for which $X$ is $\beta$-self-similar (25). The key is the following claim:
Lemma 6.
Let $X$ satisfy $\lim_{\beta\downarrow 0}\sup_{x\in X}\sum_{j=1}^{\infty}j^{\beta}|x_{j}|<\infty$. Then there exists $\beta_{0}>0$ such that $\beta\mapsto s(\beta):=\sup_{x\in X}\sum_{j=1}^{\infty}j^{\beta}|x_{j}|$ is continuous on $\left[{0},{\beta_{0}}\right)$, and $\lim_{\beta\downarrow 0}s(\beta)=\sup_{x\in X}\left\|{x}\right\|_{1}$.
Proof.
Let $\beta_{0}=\sup\{\beta>0\mid s(\beta)<\infty\}$. Take $\beta\in\left[{0},{\beta_{0}}\right)$, and let $\beta^{\prime}\to\beta$ in $\left[{0},{\beta_{0}}\right)$. By the assumption that $\sup_{x\in X}\sum_{j=1}^{\infty}j^{\beta}|x_{j}|<\infty$ for all $\beta<\beta_{0}$, for any $\epsilon>0$ and all suitably small $\gamma>0$, we may take $N$ such that $\sup_{X}\sum_{j>N}j^{\beta+\gamma}|x_{j}|<\epsilon$. Then for $\beta^{\prime}$ close enough to $\beta$, we obtain
$|s(\beta)-s(\beta^{\prime})|\leq\sup_{x\in X}\sum_{j\leq N}|j^{\beta}-j^{\beta^{\prime}}||x_{j}|+\sup_{x\in X}\sum_{j>N}(j^{\beta}+j^{\beta^{\prime}})|x_{j}|\leq\sup_{x\in X}\sum_{j\leq N}|j^{\beta}-j^{\beta^{\prime}}||x_{j}|+2\epsilon.$
As $\beta\mapsto j^{\beta}$ is continuous in $\beta$, when $\beta^{\prime}$ is close enough to $\beta$ we obtain $N|j^{\beta}-j^{\beta^{\prime}}|\leq\epsilon$ for all $j\leq N$. So $|s(\beta)-s(\beta^{\prime})|\leq\epsilon\sup_{x\in X}\left\|{x}\right\|_{1}+2\epsilon$. As $\epsilon>0$ was arbitrary this completes the proof. ∎
The preceding lemma demonstrates that $\lim_{\beta\downarrow 0}\sup_{x\in X}\sum_{j=1}^{\infty}j^{\beta}|x_{j}|=\sup_{x\in X}\left\|{x}\right\|_{1}$, so there is some $\beta>0$ for which $\sup_{x\in X}\sum_{j=1}^{\infty}j^{\beta}|x_{j}|\leq e\sup_{x\in X}\left\|{x}\right\|_{1}$. Thus $\textup{effdim}(X)<\infty$.
Finally, for the final claim of the observation, let $\mathbf{1}_{n}$ denote the linear functional $\langle\mathbf{1}_{n},x\rangle=\sum_{j=1}^{n}x_{j}$, and let $N_{\gamma}$ be the smallest $N$ such that $\sup_{x\in X}\langle\mathbf{1}_{N},x\rangle\geq\sup_{x\in X}n^{2\gamma}\sum_{j>n}|x_{j}|$ for all $n\geq N$. Define
$N:=\max\left\{N_{\gamma},\exp\left(1+\frac{3}{2\gamma}\log\frac{3}{2\gamma}\right),\exp\left(\frac{1}{2\log(e-\sqrt{e})}\right)\right\}.$
We claim that $X$ is $\beta$-self-similar (25) for $\beta=\frac{1}{2\log N}$, so that $\textup{effdim}(X)\leq\exp(1/\beta)=N^{2}$ as desired. To see the claim of self-similarity, note that the choice $\beta=\frac{1}{2\log N}$ guarantees $\beta-2\gamma\leq-\frac{3\gamma}{2}$, and
$\sum_{j=1}^{\infty}j^{\beta}|x_{j}|$ $\leq N^{\frac{1}{2\log N}}\sum_{j=1}^{N}|x_{j}|+\sum_{j=N+1}^{\infty}j^{\beta}|x_{j}|\leq\sqrt{e}\sup_{x\in X}\langle\mathbf{1}_{N},x\rangle+\sum_{j=N+1}^{\infty}j^{\beta}|x_{j}|.$
Define the index blocks $\mathcal{B}_{k}:=\{i\in\mathbf{N}\mid e^{k}\leq i<e^{k+1}\}$, so that
$\sum_{j=N+1}^{\infty}j^{\beta}|x_{j}|$ $\leq\sum_{k=\lfloor{\log N}\rfloor}^{\infty}\sum_{j\in\mathcal{B}_{k}}j^{\beta}|x_{j}|\leq\sum_{k=\lfloor{\log N}\rfloor}^{\infty}e^{\beta(k+1)}\frac{1}{e^{2k\gamma}}\sup_{x\in X}\langle\mathbf{1}_{N},x\rangle$
because $\sum_{j\in\mathcal{B}_{k}}|x_{j}|\leq e^{-2k\gamma}\sup_{x\in X}\langle\mathbf{1}_{N},x\rangle$ by assumption. Computing the infinite sums, we have that because $\frac{1}{2\log N}\leq\frac{\gamma}{2}$,
$\sum_{k=\lfloor{\log N}\rfloor}^{\infty}e^{\beta(k+1)}\frac{1}{e^{2k\gamma}}$ $=e^{\beta}\sum_{k=\lfloor{\log N}\rfloor}^{\infty}e^{k(\beta-2\gamma)}\leq e^{\beta}e^{-3\gamma\lfloor{\log N}\rfloor/2}\sum_{k=0}^{\infty}e^{-3k\gamma/2}$
$=\frac{e^{\beta-3\gamma\lfloor{\log N}\rfloor/2}}{1-e^{-3\gamma/2}}\leq\left(\frac{N}{e}\right)^{-\frac{3}{2}\gamma}\frac{2e^{\beta}}{3\gamma},$
where we used that $e^{x}\geq 1+x$, or $1-e^{x}\leq-x$. Finally, noting that $(N/e)^{-\kappa}\leq\kappa$ iff $\log N\geq 1+\frac{1}{\kappa}\log\frac{1}{\kappa}$, we substitute $\kappa=\frac{3\gamma}{2}$ to obtain $\sum_{j=N+1}^{\infty}j^{\beta}|x_{j}|\leq e^{\beta}\sup_{x\in X}\langle\mathbf{1}_{N},x\rangle$. We have shown that
$\sum_{j=1}^{\infty}j^{\beta}|x_{j}|\leq(\sqrt{e}+e^{\beta})\sup_{x\in X}\langle\mathbf{1}_{N},x\rangle$
for $\beta=\frac{1}{2\log N}$. So long as $\beta\leq\log(e-\sqrt{e})$ we have $\sqrt{e}+e^{\beta}\leq e$.
Appendix D Proof of Theorem 5
As in the proof of Theorem 4, let $A\succ 0$ be a positive definite matrix, so that the regret of (Euclidean) mirror descent with distance generating function $h_{A}(x)=\frac{1}{2}x^{\top}Ax$ is
$\mathsf{Regret}_{k,A}(x)=\bigg{\langle}\sum_{i\leq k}g_{i},x\bigg{\rangle}+\frac{1}{2}\sum_{i\leq k}\left\|{g_{i}}\right\|_{A^{-1}}^{2}-\frac{1}{2}\bigg{\|}{\sum_{i\leq k}g_{i}}\bigg{\|}_{A^{-1}}^{2}.$
For vectors $u,v\in G$ to be chosen and $p\in(0,1)$ to be chosen as well, we set
  1. (a)
    $g_{i}=u$ for $pk/2$ of the indices $i\in[k]$
  2. (b)
    $g_{i}=-u$ for $pk/2$ of the indices $i\in[k]$
  3. (c)
    $g_{i}=v$ for $(1-p)k$ of the indices $i\in[k]$.
This then yields regret lower bound
$\mathsf{Regret}_{k,A}(x)\geq(1-p)kv^{\top}x-\frac{(1-p)^{2}k^{2}}{2}v^{\top}A^{-1}v+\frac{pk}{2}u^{\top}A^{-1}u.$ (26)
We argue we may assume w.l.o.g. that $\limsup_{k}\left|\!\left|\!\left|{A(k)}\right|\!\right|\!\right|_{\rm op}/\sqrt{k}<\infty$. Suppose to the contrary that along a subsequence, which for simplicity we take to be the entire sequence, that $\left|\!\left|\!\left|{A(k)}\right|\!\right|\!\right|_{\rm op}\gg\sqrt{k}$. Let $C<\infty$ be arbitrary, take $k$ large enough that $\left|\!\left|\!\left|{A(k)}\right|\!\right|\!\right|_{\rm op}\geq C\sqrt{k}$, and assume w.l.o.g. that $C\leq\sqrt{k}$ (we can always take $k$ larger). Let $\delta_{0}>0$ be such that $\mathbf{B}_{2}(0,\delta_{0})\subset G$, which of course exists because $G$ is a convex body. Let $w$ be the unit eigenvector of $A=A(k)$ achieving $w^{\top}Aw=\left|\!\left|\!\left|{A}\right|\!\right|\!\right|_{\rm op}$, and set $v=\delta w$ for a $0\leq\delta\leq\delta_{0}$ to be chosen, so that $v=\delta w\in G$. Then at such indices $k$ the lower bound (26) implies
$\mathsf{Regret}_{k,A(k)}(x)$ $\geq(1-p)k\delta w^{\top}x-\frac{(1-p)^{2}k^{2}\delta^{2}}{2\left|\!\left|\!\left|{A(k)}\right|\!\right|\!\right|_{\rm op}}+\frac{kp}{2}u^{\top}A^{-1}u$
$\geq\frac{k(1-p)\delta}{2}\left[w^{\top}x-\frac{(1-p)\sqrt{k}\delta}{2C}\right].$
Taking a supremum over $x\in X$, which is a convex body (so that it has interior), we have $\sup_{x\in X}w^{\top}x=c(X)>0$, whence for large enough $k$ we have
$\sup_{x\in X}\mathsf{Regret}_{k,A(k)}(x)$ $\geq\sup_{0\leq\delta\leq\delta_{0}}k(1-p)\delta\left[c(X)-\frac{(1-p)\sqrt{k}\delta}{2C}\right]$
$\geq\frac{C(1-p)\cdot\min\{c(X),1\}}{2}\sqrt{k},$
where the second inequality sets $\delta=C\cdot\min\{c(X),1\}/\sqrt{k}$, which satisfies $\delta\leq\delta_{0}$ for $k$ large enough. As $c(X)>0$ and $C<\infty$ was otherwise arbitrary, whenever $p<1$ the preceding lower bound is stronger than that the theorem claims.
Returning to the main thread, we may therefore assume that $\sup_{k}\left|\!\left|\!\left|{A(k)}\right|\!\right|\!\right|_{\rm op}/\sqrt{k}\leq C$ for some finite $C$. Returning to the regret bound (26), we optimize over $v$ and $u$ to obtain
$\mathsf{Regret}_{k,A}(x)\geq\sup_{u,v\in G}\left[k(1-p)v^{\top}x-\frac{k^{2}(1-p)^{2}}{2}v^{\top}A^{-1}v+\frac{kp}{2}u^{\top}A^{-1}u\right].$
Considering the supremum over $v$, we have for any $A\succ 0$ that
$\mathop{\textup{argmax}}_{v}\left\{k(1-p)v^{\top}x-\frac{k^{2}(1-p)^{2}}{2}v^{\top}A^{-1}v\right\}=\frac{1}{k(1-p)}Ax.$
Because $X$ is bounded and $\left|\!\left|\!\left|{A}\right|\!\right|\!\right|_{\rm op}\leq C/\sqrt{k}$, for $p<1$ and suitably large $k$ the $v$ achieving this supremum evidently satisfies $\left\|{v}\right\|_{2}=\left\|{Ax}\right\|_{2}/(k(1-p))\leq\delta_{0}$, so that $v\in G$. Then for any (fixed) $p\in(0,1)$, for all large enough $k$ we obtain
$\mathsf{Regret}_{k,A}(x)\geq\sup_{u\in G}\frac{1}{2}\left[kpu^{\top}A^{-1}u+x^{\top}Ax\right].$ (27)
We use a duality argument to lower bound the quantity (27). Let $\mathcal{P}(X)$ and $\mathcal{P}(G)$ denote the collections of probability measures on $X$ and $G$, respectively. Then
$\inf_{A\succeq 0}\sup_{x\in X}\sup_{u\in G}\left\{kpu^{\top}A^{-1}u+x^{\top}Ax\right\}$
$\geq\inf_{A\succeq 0}\sup_{\nu\in\mathcal{P}(X)}\sup_{\mu\in\mathcal{P}(G)}\left[kp\langle A^{-1},\mathbf{E}_{\mu}[uu^{\top}]\rangle+\langle A,\mathbf{E}_{\nu}[xx^{\top}]\rangle\right]$
$\geq\sup_{\nu\in\mathcal{P}(X)}\sup_{\mu\in\mathcal{P}(G)}\inf_{A\succeq 0}\left\{kp\cdot\langle A^{-1},\mathbf{E}_{\mu}[uu^{\top}]\rangle+\langle A,\mathbf{E}_{\nu}[xx^{\top}]\rangle\right\}.$
For shorthand, fix $X_{\mu}=\mathbf{E}_{\nu}[xx^{\top}]$ and $G_{\mu}=\mathbf{E}_{\mu}[uu^{\top}]$. Then taking derivatives with respect to $A$, we see that the inner infimum is achieved whenever
$-kpA^{-1}G_{\mu}A^{-1}+X_{\nu}=0,~{}~{}\mbox{i.e.}~{}~{}AX_{\nu}A=kpG_{\mu}.$ (28)
For $X\succ 0$ and $G\succ 0$, the matrix equation $AXA=G$ has solution
$A=X^{-1/2}(X^{1/2}GX^{1/2})^{1/2}X^{-1/2},$
so the solution (28) satisfies
$A=\sqrt{kp}\,X_{\nu}^{-1/2}\left(X_{\nu}^{1/2}G_{\mu}X_{\nu}^{1/2}\right)^{1/2}X_{\nu}^{-1/2}.$
We must argue that (for an appropriate constant $C<\infty$) we have $\left|\!\left|\!\left|{A}\right|\!\right|\!\right|_{\rm op}\leq C\sqrt{k}$ for this choice of $A$. But for any finite $b<\infty$, so long as $\nu$ satisfies $\mathbf{E}_{\nu}[xx^{\top}]\succeq b^{-1}I_{n}$, we certainly have $\left|\!\left|\!\left|{A}\right|\!\right|\!\right|_{\rm op}\leq C\sqrt{k}$, where $C<\infty$ grows at most as $1/b$ and depends on the diameters of $X$ and $G$. Substituting this value for $A$ then yields
$\frac{1}{2}kp\cdot\langle A^{-1},G_{\mu}\rangle+\langle A,X_{\nu}\rangle=\sqrt{kp}\cdot{\rm tr}\left(\left(X_{\nu}^{1/2}G_{\mu}X_{\nu}^{1/2}\right)^{1/2}\right).$
Substituting into the regret lower bound (27), we have for any $b<\infty$ and $p\in(0,1)$ that for all large enough $k$
$\sup_{x\in X}\mathsf{Regret}_{k,A}(x)$
$\geq\sqrt{kp}\sup_{\nu\in\mathcal{P}(X)}\sup_{\mu\in\mathcal{P}(G)}\left\{{\rm tr}\left(\left(\mathbf{E}_{\nu}[xx^{\top}]^{1/2}\mathbf{E}_{\mu}[gg^{\top}]\mathbf{E}_{\nu}[xx^{\top}]^{1/2}\right)^{1/2}\right)\mid\mathbf{E}_{\nu}[xx^{\top}]\succeq b^{-1}I_{n}\right\}.$
Finally, we use the following lemma relating quadratic hulls and measures.
Lemma 7.
Let $X$ and $G$ be orthosymmetric convex bodies and $\mathcal{P}(X)$ denote the collection of probability measures on $X$ (and similarly for $G$). Define
$R(\mu,\nu):={\rm tr}\left(\left(\mathbf{E}_{\nu}[xx^{\top}]^{1/2}\mathbf{E}_{\mu}[gg^{\top}]\mathbf{E}_{\nu}[xx^{\top}]^{1/2}\right)^{1/2}\right).$
Then
$\sup_{\mu\in\mathcal{P}(G)}\sup_{\nu\in\mathcal{P}(X)}R(\mu,\nu)$ $=\sup_{\mu\in\mathcal{P}(G)}\sup_{\nu\in\mathcal{P}(X)}\sum_{j=1}^{n}\sqrt{\mathbf{E}_{\nu}[x_{j}^{2}]\mathbf{E}_{\mu}[g_{j}^{2}]}=\sup_{q\in\mathsf{QHull}(X)}\sup_{y\in\mathsf{QHull}(G)}\langle q,y\rangle.$
Moreover, the suprema can be taken over symmetric measures.
Proof.
We prove the first equality first. The function $R(\mu,\nu)$ is jointly concave in $\mu$ and $\nu$, as we may write it as the infimum $R(\mu,\nu)=\frac{1}{2}\inf_{A\succeq 0}\{\langle A^{-1},\mathbf{E}_{\mu}[gg^{\top}]\rangle+\langle A,\mathbf{E}_{\nu}[xx^{\top}]\rangle\}$ of linear functions of $\mu$ and $\nu$. Fix measures $\mu_{+}$ and $\nu_{+}$, and for a diagonal matrix of signs $S$, let $\mu_{-}$ and $\nu_{-}$ be the induced measures on $g\sim\mu$ and $x\sim\nu$. Then note that
$R(\mu_{-},\nu_{-})=\frac{1}{2}\inf_{A\succ 0}{\rm tr}\left(\langle SA^{-1}S,\mathbf{E}_{\mu_{+}}[gg^{\top}]\rangle+\langle SAS,\mathbf{E}_{\nu_{+}}[xx^{\top}]\rangle\right)=R(\mu_{+},\nu_{+}),$
because $SA^{-1}S=(SAS)^{-1}$, as $S^{2}=I$. (Joint) concavity then yields
$R\left(\frac{1}{2}(\mu_{+}+\mu_{-}),\frac{1}{2}(\nu_{+}+\nu_{-})\right)\geq\frac{1}{2}R(\mu_{+},\nu_{+})+\frac{1}{2}R(\mu_{-},\nu_{-})=R(\mu_{+},\nu_{+}).$
Then for any measure $\mu$, if we allow $\overline{\mu}$ to be the induced measure on $Sx$ for $x\sim\mu$ and $S$ drawn uniformly from diagonal sign matrices, and similarly for $\nu$ and $\overline{\nu}$, we obtain
$R\left(\overline{\mu},\overline{\nu}\right)\geq R(\mu,\nu).$
Of course, $\mathbf{E}_{\overline{\mu}}[gg^{\top}]=\mathbf{E}_{\mu}[Sgg^{\top}S]=\textup{diag}(\mathbf{E}_{\mu}[g_{j}^{2}]_{j=1}^{n})$ is diagonal, as is $\mathbf{E}_{\overline{\nu}}[xx^{\top}]$, and so
$\sup_{\mu,\nu}R(\mu,\nu)=\sup_{\mu,\nu}\sum_{j=1}^{n}\sqrt{\sqrt{\mathbf{E}_{\nu}[x_{j}^{2}]}\mathbf{E}_{\mu}[g_{j}^{2}]\sqrt{\mathbf{E}_{\nu}[x_{j}^{2}]}}=\sup_{\mu,\nu}\sum_{j=1}^{n}\sqrt{\mathbf{E}_{\mu}[g_{j}^{2}]}\sqrt{\mathbf{E}_{\nu}[x_{j}^{2}]}.$
For the second equality, recall that for any vector $q\in\mathsf{QHull}(X)$ with $q\succeq 0$, we may write $q=\sqrt{z}$ (applied elementwise), where $z$ is a convex combination of vectors of the form $[x_{j}^{2}]_{j=1}^{n}$, $x\in X$. Let $\nu\in\mathcal{P}(X)$. Then for $x^{i}\stackrel{{\scriptstyle\textup{iid}}}{{\sim}}\nu$, the strong law of large numbers guarantees that $\frac{1}{n}\sum_{i=1}^{n}x^{i}(x^{i})^{\top}\to\mathbf{E}_{\nu}[xx^{\top}]$ with probability 1, and so for any measure $\nu\in\mathcal{P}(X)$ and any $\epsilon>0$, there exists a finite set of vectors $\{x^{i}\}_{i=1}^{m}$ with $|\frac{1}{m}\sum_{i=1}^{m}(x^{i}_{j})^{2}-\mathbf{E}_{\nu}[x_{j}^{2}]|\leq\epsilon$. The same is true for $\mu\in\mathcal{P}(G)$, and so a continuity argument gives the second equality. ∎
By a slight perturbation, we therefore obtain that for any $\epsilon>0$, we can choose $b$ large enough that for all large $k$, using the $R$ in Lemma 7 we have
$\sup_{x\in X}\mathsf{Regret}_{k,A(k)}(x)$ $\geq\sqrt{kp}\sup_{\mu\in\mathcal{P}(G),\nu\in\mathcal{P}(X)}\left\{R(\mu,\nu)\mid\mathbf{E}_{\nu}[xx^{\top}]\succeq b^{-1}I_{n}\right\}$
$\geq(1-\epsilon)\sqrt{kp}\sup_{q\in\mathsf{QHull}(X)}\sup_{y\in\mathsf{QHull}(G)}\langle q,y\rangle.$
As $\epsilon>0$ and $p<1$ were arbitrary, this completes the proof of the lower bound.
The upper bound on the regret follows immediately from Corollary 7 once we recognize that any orthosymmetric compact convex set $G$ defines a norm via $\gamma(g)=\inf\{t>0\mid t^{-1}g\in G\}$ (see [24, Def. V.1.2.4 and Ch. V.3.2]).
Appendix E Proofs related to $\ell_{1}$-diameters and $n$-widths
Here, we collect the lemmas necessary to prove Corollary 12. Recall that $\mathbf{1}_{n}\in\mathbf{R}^{\mathbf{N}}$ denotes the vector with $1$ in the first $n$ positions and 0 elsewhere.
Lemma 8.
For the set $X=\textup{Conv}(C_{0}\cup C_{1})$ that equation (19) defines, we have
$\sup_{q\in\mathsf{QHull}(X)}\langle\mathbf{1}_{n},q\rangle=\sqrt{\sum_{j=1}^{n}a_{j}^{-2}}~{}~{}\mbox{and}~{}~{}\sup_{x\in X}\langle\mathbf{1}_{n},x\rangle=\max_{m\leq n}\frac{\sum_{j=1}^{m}b_{j}}{\sqrt{\sum_{j=1}^{m}a_{j}^{2}b_{j}^{2}}}=\max_{m\leq n}\frac{\langle\mathbf{1}_{m},b\rangle}{Z(m)}.$
Proof.
Let $Q=\mathsf{QHull}(X)=\{q\mid\sum_{j=1}^{\infty}a_{j}^{2}q_{j}^{2}\leq 1\}$. By Cauchy-Schwarz, the suprema of $\langle\mathbf{1}_{n},q\rangle$ over $q\in Q$ satisfy $q_{j}=\frac{\lambda}{a_{j}^{2}}$ where $\lambda>0$ normalizes $q$ so that $\sum_{j=1}^{n}a_{j}^{2}q_{j}^{2}=1$, that is, $\lambda=(\sum_{j=1}^{n}a_{j}^{-2})^{-1/2}$. For the second equality, note that $\langle\mathbf{1}_{n},x\rangle$ is linear in $x$, and so the supremum is achieved at one of the vertices of $C_{0}$ or $C_{1}$. Thus
$\sup_{x\in X}\langle\mathbf{1}_{n},x\rangle=\max_{x\in C_{0}}\langle\mathbf{1}_{n},x\rangle\vee\max_{x\in C_{1}}\langle\mathbf{1}_{n},x\rangle=\max_{j\leq n}\frac{1}{a_{j}}\vee\max_{m\leq n}\frac{1}{Z(m)}\langle\mathbf{1}_{m},b\rangle.$
Substitute $1=\max_{j\leq n}\frac{1}{a_{j}}$, as $a_{1}=1$ and $a_{j}$ are nondecreasing, then recognize that $b_{1}/\sqrt{b_{1}^{2}}=1$ to obtain the lemma. ∎
We now give rough bounds on the widths of the set $X$ and its hull.
Lemma 9.
For the set $X=\textup{Conv}(C_{0}\cup C_{1})$, we have
$0pt^{2}(n)=\sup_{m\geq n}\frac{m-n}{\sum_{j=1}^{m}a_{j}^{2}}$
and
$0pt^{2}(n)\geq w_{\textup{nl}}^{2}(n)\geq\sup_{m\geq n}\frac{1}{Z(m)^{2}}\sum_{j=n+1}^{m}b_{j}^{2}=\sup_{m\geq n}\frac{\sum_{j=n+1}^{m}b_{j}^{2}}{\sum_{j=1}^{m}a_{j}^{2}b_{j}^{2}}.$
Proof.
For the linear width, we recognize that $Q=\{q\mid\sum_{j=1}^{\infty}a_{j}^{2}q_{j}^{2}\leq 1\}$ is elliptical, so using the characterization (18) of $0pt^{2}(n)$ gives the first claim of the lemma. For the second we can take as a lower bound the nonlinear width of the set $C_{1}$, so that
$w_{\textup{nl}}^{2}(n)\geq\sup_{x\in C_{1}}\bigg{\{}\sum_{j>n}x_{(j)}^{2}\bigg{\}}=\sup_{m\geq n}\bigg{\{}\frac{1}{Z(m)^{2}}\sum_{j=n+1}^{m}b_{j}^{2}\bigg{\}}$
as desired. ∎
Finally, we take the scalars $a_{j}$, $b_{j}$ as in the statement of Corollary 12. Set
$a_{j}=j^{\alpha/2}~{}~{}\mbox{and}~{}~{}b_{j}^{2}=2^{j},$
where $10<\alpha<$ 1. Then direct calculations yield the asymptotics that for $m\geq n$,
[…]
$2^{m}m^{\alpha}=a_{m}^{2}b_{m}^{2}\leq\sum_{j=1}^{m}a_{j}^{2}b_{j}^{2}$ $=2^{m}m^{\alpha}\sum_{j=1}^{m}\left(\frac{j}{m}\right)^{\alpha}2^{j-m}\leq 2^{m}m^{\alpha}\sum_{j=0}^{m-1}2^{-j}\leq 2^{m+1}m^{\alpha}$
$2^{m}=b_{m}^{2}\leq\sum_{j=n+1}^{m}b_{j}^{2}$ $\leq\sum_{j=1}^{m}2^{j}\leq 2^{m+1}.$ (29)
The first equation in (29) implies that
$0pt^{2}(n)=\sup_{m\geq n}\frac{m-n}{\sum_{j=1}^{m}a_{j}^{2}}\asymp\sup_{m\geq n}\frac{m-n}{m^{1+\alpha}}\asymp n^{-\alpha},$
while the last two equations lower bound the nonlinear width (via Lemma 9) by
$w_{\textup{nl}}^{2}(n)\geq\sup_{m\geq n}\frac{\sum_{j=n+1}^{m}b_{j}^{2}}{\sum_{j=1}^{m}a_{j}^{2}b_{j}^{2}}\geq\sup_{m\geq n}\frac{2^{m}}{2^{m}m^{\alpha}}=n^{-\alpha},$
and so we have $0pt^{2}(n)\lesssim w_{\textup{nl}}^{2}(n)\leq 0pt^{2}(n)$ for all $n$.
To prove Corollary 12, it remains to compute $\ell_{1}$ diameter ratio. Applying Lemma 8, for $\alpha<1$ we have
$\sup_{q\in\mathsf{QHull}(X)}\langle\mathbf{1}_{n},q\rangle=\sqrt{\sum_{j=1}^{n}a_{j}^{-2}}=\sqrt{\sum_{j=1}^{n}j^{-\alpha}}\asymp n^{\frac{1-\alpha}{2}}.$
On the other hand, because $\sum_{j=1}^{m}b_{j}\asymp 2^{m/2}$, the bounds (29) and Lemma 8 give
$\sup_{x\in X}\langle\mathbf{1}_{n},x\rangle=\max_{m\leq n}\frac{\sum_{j=1}^{m}b_{j}}{\sqrt{\sum_{j=1}^{m}a_{j}^{2}b_{j}^{2}}}\lesssim\max_{m\leq n}\frac{2^{m/2}}{m^{\alpha/2}2^{m/2}}=1.$